### v2

# COOL-LAMPS. VII. Quantifying Strong-lens Scaling Relations with 177 Cluster-scale Gravitational Lenses in DECaLS

Abstract
We compute parametric measurements of the Einstein-radius-enclosed total mass for 177 cluster-scale strong gravitational lenses identified by the ChicagO Optically-selected Lenses Located At the Margins of Public Surveys (COOL-LAMPS) collaboration with lens redshifts ranging from $\sim 0.2$ to $\sim 1.0$ using only two parameters: the Einstein radius and the brightest-cluster-galaxy (BCG) redshift. We constrain the Einstein-radius-enclosed luminosity and stellar mass by fitting parametric spectral energy distributions (SEDs) with aperture photometry from the Dark Energy Camera Legacy Survey (DECaLS) in the $grz$ Dark Energy Camera (DECam) filters. We find that the BCG redshift, enclosed total mass, and enclosed luminosity are strongly correlated and well described by a planar relationship in 3D space. We also find that the enclosed total mass and stellar mass are correlated with a logarithmic slope of $0.443\pm 0.035$, and the enclosed total mass and stellar-to-total mass fraction are correlated with a logarithmic slope of $-0.563\pm 0.035$. The correlations described here can be used to validate strong lensing candidates in upcoming imaging surveys—such as Rubin/Legacy Survey of Space and Time (LSST)—in which an algorithmic treatment of lensing systems will be needed due to the sheer volume of data these surveys will produce.
Galaxy clusters (584) — High-redshift galaxy clusters (2007) — Scaling relations (2031) — Spectral energy distribution (2129) — Strong gravitational lensing (1643)
[…]
1 Introduction
Strong gravitational lensing is a rare phenomenon in the universe in which an intervening object with sufficient surface mass density bends light from a background source relative to the observer such that multiple images of the source are formed. For cases in which the background source is a galaxy, the lensing effect can create highly distorted and magnified images of the source galaxy in an arclike shape (e.g., Lynds & Petrosian 1986; Soucail et al. 1987; Kneib & Natarajan 2011; Meneghetti et al. 2013; Rivera-Thorsen et al. 2017; Shajib et al. 2022; Zhang et al. 2023). The discovery of new gravitational lenses of this type is accelerating rapidly, particularly with the help of machine learning applied to finding galaxy-galaxy strong lenses in large optical-imaging surveys (e.g., Huang et al. 2021; Rojas et al. 2022; Zaborowski et al. 2022). Coupled with access to a flood of new data from large surveys such as Rubin/LSST in the coming years, the population of candidate lenses both in the bulk and at the margins will only grow. Contemporary methodologies often rely on visual inspection in order to winnow candidate lists to an acceptable number worthy of detailed follow-up (Diehl et al., 2017); such inspection has been successful in locating new systems thus far (Rojas et al., 2023). However, with thousands to millions of lines of sight to choose from in an ever-growing plethora of imaging data, statistical measures and correlations in existing lensing systems may help hone candidate-lens samples by quantifying instances of strong lensing in tandem with traditional morphological identification. This would produce a purer sample and help streamline the ability to identify targets. In this paper, we analyze a sample of 177 strong gravitational lenses identified via visual inspection by the COOL-LAMPS collaboration in DECaLS Legacy Survey Data Release 8 (LS DR8, Dey et al. 2019) images, and we investigate the scaling relations between the BCG redshift, Einstein-radius-enclosed (enclosed) total mass, enclosed luminosity, enclosed stellar mass, and enclosed stellar-to-total mass fraction therein. The sample we use here consists of a refined subsample of all strong-lensing candidates identified by the COOL-LAMPS team, and we consider them unambiguous and robust as informed by their visual morphology—despite a lack of definite spectroscopic confirmation.
Throughout this paper, we adopt a $\mathrm{\Lambda CDM}$ WMAP-9 cosmology (Hinshaw et al., 2013). All photometric calculations were done in the AB magnitude system.
2 Methodology
Galaxies and galaxy clusters host their mass in a variety of forms. Theory and observations both place strong constraints on observed stellar luminosity as a function of mass (e.g., Kuiper 1938; Presotto et al. 2014; Wang & Zhong 2018), hot intracluster gas contributes significantly to the total mass in galaxy clusters (e.g., Kellogg et al. 1972; Forman et al. 1972; Markevitch & Vikhlinin 2007), and both galaxies and clusters are composed of a significant fraction of dark matter (e.g., Zwicky 1933; Rubin 1986; Navarro et al. 1996; Bryan & Norman 1998; Hradecky et al. 2000). However, robustly measuring these mass components on a per-system basis is time and resource intensive—which is unsustainable for large samples. Conveniently, the existence of lensing prescribes a specific amount of mass that must exist to create the observed lensing in the first place.
Remolina González et al. (2020) have shown that knowing only the angle between the BCG and lensed source arc (a proxy for Einstein radius; $\theta_{E}$) and the line-of-sight geometry (redshift of the lens and the source; $z_{L}$, $z_{S}$), the core mass interior to the Einstein radius of a galaxy cluster may be measured with minimal scatter on order of $\approx 10\%$. With these three variables, the mass can be immediately calculated and accounts for all of the projected mass density interior to the Einstein radius. Measuring mass in this way also offers a well-defined aperture with which to constrain photometric properties as well. By directly integrating over pixels inside this aperture in flux-calibrated images, the photometry can be automatically constrained and utilized by an SED modeller to obtain measurements of enclosed luminosity and stellar mass.
The COOL-LAMPS collaboration has constructed a large dataset of strong lensing candidates using a simple human-led ranking system (e.g., Khullar et al. 2021). From this, we have collated a sample of 177 galaxy clusters and groups that exhibit robust visual evidence of strong gravitational lensing in which the primary lensed arc is roughly circular with its projected center located at or very near to the BCG. Any systems with complex source-galaxy morphology are less likely to be well-described by the simple symmetric dark-matter profiles that inform the parametric mass estimates described in Remolina González et al. (2020). Such complex systems were intentionally excluded from the sample considered in this work.
Figure 1: A subset of nine strong gravitational lenses used in this work. The lenses are arranged by increasing redshift from the top to the bottom of the figure, and they are arranged by increasing Einstein radius from the left to the right of the figure. Red circles denote positional constraints used to derive the Einstein radius in each of the systems. White circles represent the Einstein aperture for each system. A black ‘x’ at the center of each subplot marks the exact coordinates of the BCG. The lens redshifts and Einstein radii for each system are also shown in each panel. RGB images have equal astrometric scaling and are sourced from DECaLS LS DR9 (Dey et al., 2019) using the $z$, $r$, and $g$ bands with custom color scaling respectively.
Figure 2: Observed-frame spectral energy distributions (in maggies) for the nine systems shown in Figure 1 after SED fitting with Prospector. The top plot of each subplot contains the best-fit spectrum (red line), 1024 spectra drawn from their respective posterior distribution (grey lines), observed photometry (blue points), and photometric filter transmission curves for the Dark Energy Camera (DECam, Flaugher et al. 2015) $g$, $r$, and $z$ filters from left to right respectively (in arbitrary units). The bottom plot of each subplot shows the residual between the best-fit spectra and the observed photometry normalized by the standard deviation of the observed photometry.
2.1 Einstein Radius
We defined the Einstein radius for each lensing system as the radius of a circle centered on the BCG (Remolina González et al., 2020) which minimizes the total angular separation between each of the three most-readily-identifiable bright “clumps” in the tangentially-lensed arcs and the perimeter of said circle. Henceforth, we refer to this circular aperture centered on the BCG with radius equal to the Einstein radius as the “Einstein aperture”. Four points (the BCG center and three points along each arc) physically constrain the Einstein aperture, and we obtained them by tagging the point of peak surface brightness for the BCG and similar peaks in surface brightness in the clumps of each tangentially-lensed arc respectively. If three distinct peaks were unable to be found along the tangential arcs, the main arc was simply traced. We bootstrapped the three tangential-arc positional constraints 1000 times, and we adopted the mean and standard deviation for the resulting distribution of Einstein radii as the Einstein radius and its error respectively for each system. A descriptive visualization sampling the range of Einstein radii and lens redshifts considered in this work is shown in Figure 1.
2.2 Lens Redshift
BCGs are sui generis red-sequence cluster galaxies which can also be used as proxies to infer properties of the cluster-scale dark matter halo (e.g., Oegerle & Hoessel 1991; Lauer et al. 2014). Since we assume that the lensing systems in this work can be well-described by a symmetric dark-matter profile, we take the redshift of the BCG as interchangeable with the systematic redshift of the cluster-scale halo. The lens redshift converts the measured Einstein radius to a proper transverse distance via the angular diameter distance, and we query the Sloan Digital Sky Survey Data Release 15 (SDSS DR15, Aguado et al. 2019; Kollmeier et al. 2019) in tandem with the Legacy Survey Data Release 9 (LS DR9, Dey et al. 2019; Duncan 2022) for each system in this work to obtain BCG redshifts. If the centroid coordinates for each BCG corresponded to a warning-free spectroscopic redshift in SDSS DR15, we adopted that redshift as the lens redshift for that system along with its associated error. If a given BCG lacked any corresponding value in SDSS DR15, as is the case for the higher redshift clusters in this work and/or clusters located in SDSS-non-imaged areas, we adopted the LS DR9 photometric redshift (Zhou et al., 2023) as the lens redshift and its associated error instead. Summarized in Table 2, we obtained spectroscopic redshifts for 41 systems, and photometric redshifts for the remaining 136.
Table 1: Free Parameters used in SED fitting with Prospector
Parameter Description Priors
$\mathrm{log}(M_{tot}/M_{\odot})$ Total stellar mass formed in dex solar masses. Top Hat: [8.0, 14.0]
$\mathrm{log}(Z/Z_{\odot})$ Stellar metallicity in dex solar metallicity. Top Hat: [-1.0, 0.2]
$\lambda_{2}$ Diffuse dust optical depth. Top Hat: [0.0, 2.0]
$t_{\mathrm{age}}$ Age of the cluster in Gyr. Top Hat: [$t_{z=20}$, $t_{z_{BCG}}$]
$\tau$ SFH e-folding time in Gyr. Top Hat: [0.1, 10]
2.3 Source Redshift
In strong lensing, the source redshift is typically constrained either with spectroscopic follow-up (e.g., Sharon et al. 2020) or by inferring a photometric redshift (e.g., Cerny et al. 2018). However, obtaining these redshifts for all identified source galaxies requires prolonged additional inquiry in tension with the appeal of a robust and efficient mass estimator. Remolina González et al. (2020) have shown that substituting a single known source-galaxy redshift with a distribution of source-galaxy redshifts introduces a statistically insignificant uncertainty into the final distribution of mass measurements when compared to the magnitude of other systematic uncertainties. This distribution of lensed source redshifts is measurable (Bayliss et al., 2011a, b; Tran et al., 2022), and we simply adopt a well-described Gaussian with $\mu=2$ and $\sigma=0.2$ from Bayliss et al. (2011a) in keeping with the methodology of Remolina González et al. (2020) as the distribution of source redshifts across the entire sample.
2.4 Photometry
[…]
3 Analysis
3.1 Total Mass
Following Remolina González et al. (2020), the enclosed total mass within the Einstein aperture was calculated using Equations (1) and (2), where $D(z)$ refers to the angular diameter distance at redshift $z$, via a Monte Carlo approach. The enclosed total mass was computed 1000 times in which each calculation used a different Einstein radius, source redshift, and lens redshift randomly drawn from a Gaussian distribution with standard deviation equal to the uncertainty on each measurement and mean equal to the value for each measurement. Also described in Remolina González et al. (2020) is the need to apply an empirical correction described in Equation (3) based on the “completeness” of the lensed galaxy arcs in each system, where $f(\theta_{E})$ is a cubic polynomial function specified in Table 1 of Remolina González et al. (2020). This is because the systematic bias and scatter of $M_{\Sigma}(<\theta_{E})$ has a dependence on the degree of symmetry for each system in which higher Einstein-radius systems have a stronger offset. The correction factor aims to remove this dependence as a function of Einstein radius in non-perfectly-symmetric systems.
$\Sigma_{\mathrm{cr}}(z_{L},z_{S})=\frac{c^{2}}{4\pi G}\frac{D_{S}(z_{S})}{D_{L}(z_{L})D_{LS}(z_{L},z_{S})}$ (1)
$M(<\theta_{E})=\pi(D_{L}(z_{L})\theta_{E})^{2}\,\Sigma_{\mathrm{cr}}(z_{L},z_{S})$ (2)
$\mathrm{\ }M_{\Sigma}(<\theta_{E})=\frac{M(<\theta_{E})}{f(\theta_{E})}$ (3)
For the purposes of applying this correction in each of the 177 lensing systems in this work, we measured the azimuthal coverage ($\phi$) by determining what percentage of the Einstein aperture was traced out by the lensed galaxy arcs in a given system. For example, a lens with a tangential arc stretching from an azimuthal angle of 10 to 100 degrees would have $\phi=0.25$. This was measured by determining what azimuthal fraction of the perimeter of the Einstein aperture was subtended by the the regions masking the source arcs described in Section 2.4. As was shown by Remolina González et al. (2020), whether or not the correction factor was needed was determined by using $\phi$ as an observable proxy for the degree of symmetry for each system. Larger-Einstein-radius systems with a predominantly low $\phi$ tend to deviate from a symmetric dark-matter halo and thus require the empirical correction. For spherically-symmetric systems with a large $\phi$, the measured $M_{\Sigma}(<\theta_{E})$ is taken to be fairly unbiased, and thus the correction factor is not needed. Adopting the convention used by Remolina González et al. 2020 for each system, if $\phi<0.5$, Equation (3) was applied. If $\phi\geq 0.5$, it was not applied. After accounting for the empirical correction, the mean and standard deviation of the resulting 1000 measurements of $M_{\Sigma}(<\theta_{E})$ were adopted as the enclosed total mass and error respectively for each system.
3.2 Stellar Luminosity and Mass
We conducted SED modelling using the photometry obtained in Section 2.4 with Prospector (Johnson et al., 2021; Conroy et al., 2009; Conroy & Gunn, 2010; Foreman-Mackey et al., 2013). We fit a parametric SED with three assumed parameters and five free parameters. We assumed a Chabrier initial mass function (IMF) (Chabrier, 2003), Calzetti attenuation curve for dust (Calzetti et al., 2000), and delayed-tau star-formation history (SFH) model (Carnall et al., 2019). The five free parameters that were optimized are summarized in Table 1. We utilized emcee as implemented in Prospector with 84 walkers for a total of 6720 iterations, taking only the last 840 iterations for each of the 84 walkers (70,560 total parameter vectors) as representing the posterior distribution for each of the variables of interest in order to eliminate the significant burn-in sequence of the fit. After fitting, we adopted the 50th percentile of the posterior distribution as the value for each of the five parameters, and the greater difference between the 84th-50th percentile and 50th-16th percentile values as the error. Observed-frame spectra for the systems shown in Figure 1 after fitting are shown in Figure 2.
To derive luminosity, we drew random sample of 1024 chains from the obtained posterior distribution and generated an SED for each. We integrated over the rest-frame wavelength range of 3000Å-7000Å, since this corresponds to the longest wavelength range sampled by part of at least two bands of photometry across all BCG redshifts in this work (the lower bound of 3000Å gets redshifted out of the $\mathit{g}$ and into the $\mathit{r}$ band at high redshifts). After integrating a given SED to obtain flux, we converted the observed flux to a luminosity in solar luminosities using the appropriate luminosity distance and adopted the mean and standard deviation from the 1024 random draws as the luminosity and its and error respectively for each system.
4 Results
As Figure 3 shows, the BCG redshift, enclosed total mass, and enclosed luminosity are correlated. Indeed, we note that the most massive clusters in this work contain a large amount of stars and thus also have a large luminosity. Using emcee, we fit a plane-of-best-fit using likelihood maximization taking into account the uncertainty in variables on all three axes to describe this correlation where X = $z_{{}_{BCG}}$, Y = $\mathrm{log}_{10}[M_{\Sigma}(<\theta_{E})/M_{\odot}]$, and Z = $\mathrm{log}_{10}[L(<\theta_{E})/L_{\odot}]$, yielding:
$Z=0.137_{-0.051}^{+0.051}X+0.522_{-0.023}^{+0.022}Y+4.207_{-0.290}^{+0.300}$ (4)
This simple and analyic relationship derived from a sample of known strong lenses, considering only quantities derived from observations, may have utility in differentiating bona fide strong lenses from non-lenses in large samples of candidate lenses where the evidence of lensing is not nearly as robust as the sample of clusters analyzed here. Moreover, this correlation offers a future opportunity to test simulations against observations, particularly those that aim to generate realistic galaxy populations in dense cluster environments.
Figure 3: A 3D visualization of the BCG redshift, enclosed total mass, and enclosed luminosity for the systems in this work plotted as black circles with a plane of best fit shown in black as well. Uncertainties are represented as red, blue, and green lines respectively.
A full accounting of all measured quantities in this work can be found in Table 2.
5 Discussion
5.1 Scaling Relations
Besides the enclosed total mass and luminosity, we can also infer the current enclosed stellar mass at the redshift of observation for each system using the posterior distribution of each SED fit generated with Prospector. This was obtained by correcting the derived enclosed stellar mass in each system by the surviving mass fraction returned by Prospector, since the amount of stellar mass that we see at the time of observation is less than the total stellar mass formed over the lifetime of the galaxy (e.g., Li et al. 2017). While we note that enclosed stellar mass is more complex to infer and subject to more significant systematics when compared to the more direct measures of enclosed total mass from strong lensing geometry and luminosity, it nevertheless offers another point of comparison with those measurables.
5.1.1 Total Mass — Stellar Mass
Figure 4: The total mass contained within the Einstein aperture for each system plotted against both the enclosed stellar mass and the enclosed stellar-to-total mass fraction. Points in panels I and III are color-coded according to the luminosity of the system in units of solar luminosities from stellar light within the Einstein aperture as measured in the rest-frame wavelength range from 3000Å to 7000Å. Points in panels II and IV only represent those systems in which multiple cluster members fall within the Einstein aperture. Linear regression lines in all four subpanels are drawn from the posterior distribution of their respective MCMC.
In Panel I of Figure 4, we see a clearly positive correlation between the enclosed stellar mass and the enclosed total mass for each system. By performing a linear regression with emcee using a likelihood maximization that accounts for uncertainties in variables on both axes, we find that the slope of this correlation is $0.443\pm 0.035$ from the posterior distribution of the MCMC fit. We colored the points according to their luminosity to highlight how the intrinsic differences between individual systems contribute to the overall systematic scatter. In Table 2 of Kravtsov et al. (2018), they found the stellar mass of solely the BCG and $M_{500}$ (White, 2001) to be correlated with a logarithmic slope of $0.33\pm 0.11$. In addition, Kravtsov et al. (2018) found the stellar mass of the entire cluster and $M_{500}$ to be correlated with a logarithmic slope of $0.59\pm 0.08$. The fact that the slope for this work falls between the two measurements quoted by Kravtsov et al. (2018) is consistent with how our methodology samples a physical regime between exclusively the BCG and the entire cluster.
Panel II of Figure 4 further explores this correlation, with only those clusters where multiple cluster members were contained within the Einstein aperture are plotted. A linear regression for the multiple-galaxy systems gives a slope of $0.453\pm 0.049$, which is consistent with the slope using all systems.
5.1.2 Total Mass — Stellar-to-Total Mass Fraction
In Panel III of Figure 4, we compare the enclosed stellar-to-total mass fraction against the enclosed total mass and see a clearly negative correlation (again with the same colormapping as Panel I to emphasize intrinsic scatter). We find that the slope of this correlation is $-0.563\pm 0.035$. Andreon (2010) measured 52 clusters over a slightly larger mass range and found that the stellar mass fraction depended on halo mass with a slope of $-0.55\pm 0.08$ when measuring to a physical extent of $R_{200}$. Categorizing in Panel IV identically to Panel II, we find that the multiple-galaxy systems are correlated with a slope of $-0.561\pm 0.050$. We again find that the multiple-galaxy systems are statistically indistinguishable from the slope for the entire sample shown in Panel III and in Andreon (2010). While the linear fit derived by Andreon (2010) has a slightly larger normalization, this is to be expected because the cluster masses are constrained out to $R_{200}$.
5.2 Photometric Bias and Error
Deriving light via directly integrating over the Einstein aperture presents some challenges. Since we identify red-sequence cluster members by using observed color as a proxy for cluster membership, it is possible that we may have included intrinsically-bluer galaxies at a higher redshift that appear to be the same color as the cluster members. Conversely, we may have excluded intrinsically-bluer bona fide cluster members. However, star-forming cluster members have less stellar mass and are intrinsically rarer in cluster cores (e.g., Dressler et al. 1997, 2004), and we do not expect that these missed cluster galaxies introduce a significant loss to the measured cluster light. In addition, bluer galaxies at a higher redshift appear intrinsically fainter, which are also unlikely to bias our methodology of measuring cluster light. Interlopers such as foreground stars that appear in front of cluster galaxies are forced to be masked, but any underlying galaxy flux is typically minimal due to the limited angular size of such stars.
In general, we expect that all of the unmasked flux from cluster cores at the angular scales considered in this work originates from genuine cluster members. Any modest background correction that could be made will also be effectively captured in the redshift dependence of the fits (e.g., Figure 3). While we must also consider associated galaxy populations in the immediate large-scale structure when considering galaxy clusters projected onto the sky—as such structures may also be projected onto the cluster core—those galaxies will also contribute to the lensing effect and will have their mass and light effectively captured by the methodology presented here.
There are also extremely few examples of systems whose Einstein radius is less than $\sim 2^{\prime\prime}$. To some extent, this is a consequence of the chosen sample of galaxy clusters and groups, which we would generally expect to have larger Einstein radii. Systems whose arcs are separated from the central galaxy at small radii may be located directly within the discernable BCG light. This would not only make them hard to visually notice, but it would also make them unsuitable for this analysis because the source light would be significantly blended together with the lensing cluster light. However, we caution that there is a large scatter between the total halo mass and Einstein radius (Fox et al., 2022). While we thus expect this work to undersample the small-Einstein-radius regime, it is not apparent that this significantly alters any of our results.
In addition, our use of relatively shallow ground-based imaging limits the $5\sigma$ point-source magnitude depth of objects in the $grz$ bands to $\sim$ 24.7, 23.9, and 23.0 respectively in the average case of two exposures per filter (Dey et al., 2019). Clusters higher than redshift $\sim 1$ are barely visible even in $\mathit{z}$-band filter images in the current data and are simply a rarer type of lensing system in the universe (Li et al., 2019). Both the data depth and the intrinsic redshift distribution of lenses shape the redshift distribution in this work.
5.3 Alternative Mass Measurement
In constraining mass, we also considered a second method for measuring the cluster-centric mass elucidated in Remolina González et al. (2021a) using the parametric lens modelling software LENSTOOL (Jullo et al., 2007) to measure the enclosed mass as opposed to the simple evaluation of Equations (1), (2), and (3). For the first 35 clusters that were analyzed, we generated single-halo lens models by using the three coordinates corresponding to the locations of the brightest clumps in each lensed arc as positional constraints with a single pseudo-isothermal elliptical mass distribution (PIEMD, Kassiola & Kovner 1993) locked to the center of the BCG. Mass estimates were then obtained from the resulting best-fit models according to the methodology described in Remolina González et al. (2021a).
Figure 5: A comparison of the single-halo LENSTOOL generated mass estimates ($M_{SHM}$) and the parametrically generated mass estimates ($M_{\Sigma}$) for the first 35 clusters analyzed in this work. The black line represents an ideal 1-1 correspondence, and the blue lines represent the uncertainty on the real correspondence.
As expected from Remolina González et al. (2021b), the two different methodologies show similar results, and the two mass measurements for the 35 clusters studied here are correlated with a slope of $0.960\pm 0.036$ and intercept of $0.492\pm 0.476$ as shown in Figure 5. We abandoned the LENSTOOL approach throughout the rest of the work, i.e. we find no evidence for a significant difference between these two methods.
In theory, lens modelling has the potential to be the most accurate mass-constraining methodology. However, this requires high-spatial-resolution imaging data and ample time. The process of iteratively refining just one lens model may take upwards of an hour for even a single relatively-simple system, which is untenable when the number of candidate systems reaches into the thousands. Moreover, the accuracy of measuring mass with lens modelling is sensitive to the choice of defining multiple image families along the lensed arc in a way that simple circle fitting is not. These essential positional constraints are often unresolved with even the highest-quality ground-based data. Conversely, the positional constraints that inform the Einstein radius ultimately used by Equations (1), (2), and (3) need not be extremely accurate with respect to exactly constraining image family positions. As long as the positional constraints are reasonably placed within the arcs, something which is resolved by ground-based data, it is accurate to within the systematic uncertainty.
6 Summary and Conclusions
This paper presents measurements of mass and light in 177 strong-lensing systems at cluster and group mass scales. By directly obtaining the Einstein radius and BCG redshift for each system, we compute the Einstein-radius-enclosed total mass parametrically. We use the enclosed $grz$ photometry and Prospector to parameterize the enclosed luminosity and stellar mass, and we obtain a planar relationship between the BCG redshift, enclosed total mass, and enclosed luminosity of the form $Z=0.137_{-0.051}^{+0.051}X+0.522_{-0.023}^{+0.022}Y+4.207_{-0.290}^{+0.300}$ where X = $z_{{}_{BCG}}$, Y = $\mathrm{log}_{10}[M_{\Sigma}(<\theta_{E})/M_{\odot}]$, and Z = $\mathrm{log}_{10}[L(<\theta_{E})/L_{\odot}]$. We find that the enclosed total mass and stellar mass for the entire sample of strong lensing systems considered in this work are correlated with a logarithmic slope, normalization, and scatter of $0.443\pm 0.035$, $5.958\pm 0.454$, and $0.173\pm 0.013$ respectively. We also find that the enclosed total mass and stellar-to-total mass fraction are correlated with a logarithmic slope, normalization, and scatter of $-0.563\pm 0.035$, $6.034\pm 0.457$, and $0.161\pm 0.013$ respectively. Enormous volumes of imaging data from new space-based and ground-based surveys at deeper and sharper photometric limits are on the horizon. Probing these upcoming datasets with both visual and machine-learning driven searches for new strong lenses at a variety of scales is inevitable. The correlations shown in this work should have utility in filtering rankings of possible strong lensing systems as well as offering another measure of cluster cores that can be compared to cosmological simulations.
Acknowledgments
This work is supported by The College undergraduate program and the College Innovation Fund at the University of Chicago and the Department of Astronomy and Astrophysics at the University of Chicago.
The Legacy Surveys consist of three individual and complementary projects: the Dark Energy Camera Legacy Survey (DECaLS; Proposal ID #2014B-0404; PIs: David Schlegel and Arjun Dey), the Beijing-Arizona Sky Survey (BASS; NOAO Prop. ID #2015A-0801; PIs: Zhou Xu and Xiaohui Fan), and the Mayall z-band Legacy Survey (MzLS; Prop. ID #2016A-0453; PI: Arjun Dey). DECaLS, BASS and MzLS together include data obtained, respectively, at the Blanco telescope, Cerro Tololo Inter-American Observatory, NSF’s NOIRLab; the Bok telescope, Steward Observatory, University of Arizona; and the Mayall telescope, Kitt Peak National Observatory, NOIRLab. Pipeline processing and analyses of the data were supported by NOIRLab and the Lawrence Berkeley National Laboratory (LBNL). The Legacy Surveys project is honored to be permitted to conduct astronomical research on Iolkam Du’ag (Kitt Peak), a mountain with particular significance to the Tohono O’odham Nation.
NOIRLab is operated by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation. LBNL is managed by the Regents of the University of California under contract to the U.S. Department of Energy.
This project used data obtained with the Dark Energy Camera (DECam), which was constructed by the Dark Energy Survey (DES) collaboration. Funding for the DES Projects has been provided by the U.S. Department of Energy, the U.S. National Science Foundation, the Ministry of Science and Education of Spain, the Science and Technology Facilities Council of the United Kingdom, the Higher Education Funding Council for England, the National Center for Supercomputing Applications at the University of Illinois at Urbana-Champaign, the Kavli Institute of Cosmological Physics at the University of Chicago, Center for Cosmology and Astro-Particle Physics at the Ohio State University, the Mitchell Institute for Fundamental Physics and Astronomy at Texas A&M University, Financiadora de Estudos e Projetos, Fundacao Carlos Chagas Filho de Amparo, Financiadora de Estudos e Projetos, Fundacao Carlos Chagas Filho de Amparo a Pesquisa do Estado do Rio de Janeiro, Conselho Nacional de Desenvolvimento Cientifico e Tecnologico and the Ministerio da Ciencia, Tecnologia e Inovacao, the Deutsche Forschungsgemeinschaft and the Collaborating Institutions in the Dark Energy Survey. The Collaborating Institutions are Argonne National Laboratory, the University of California at Santa Cruz, the University of Cambridge, Centro de Investigaciones Energeticas, Medioambientales y Tecnologicas-Madrid, the University of Chicago, University College London, the DES-Brazil Consortium, the University of Edinburgh, the Eidgenossische Technische Hochschule (ETH) Zurich, Fermi National Accelerator Laboratory, the University of Illinois at Urbana-Champaign, the Institut de Ciencies de l’Espai (IEEC/CSIC), the Institut de Fisica d’Altes Energies, Lawrence Berkeley National Laboratory, the Ludwig Maximilians Universitat Munchen and the associated Excellence Cluster Universe, the University of Michigan, NSF’s NOIRLab, the University of Nottingham, the Ohio State University, the University of Pennsylvania, the University of Portsmouth, SLAC National Accelerator Laboratory, Stanford University, the University of Sussex, and Texas A&M University.
BASS is a key project of the Telescope Access Program (TAP), which has been funded by the National Astronomical Observatories of China, the Chinese Academy of Sciences (the Strategic Priority Research Program “The Emergence of Cosmological Structures” Grant # XDB09000000), and the Special Fund for Astronomy from the Ministry of Finance. The BASS is also supported by the External Cooperation Program of Chinese Academy of Sciences (Grant # 114A11KYSB20160057), and Chinese National Natural Science Foundation (Grant # 12120101003, # 11433005).
The Legacy Survey team makes use of data products from the Near-Earth Object Wide-field Infrared Survey Explorer (NEOWISE), which is a project of the Jet Propulsion Laboratory/California Institute of Technology. NEOWISE is funded by the National Aeronautics and Space Administration.
The Legacy Surveys imaging of the DESI footprint is supported by the Director, Office of Science, Office of High Energy Physics of the U.S. Department of Energy under Contract No. DE-AC02-05CH1123, by the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility under the same contract; and by the U.S. National Science Foundation, Division of Astronomical Sciences under Contract No. AST-0950945 to NOAO.
The Photometric Redshifts for the Legacy Surveys (PRLS) catalog used in this paper was produced thanks to funding from the U.S. Department of Energy Office of Science, Office of High Energy Physics via grant DE-SC0007914.
Funding for the Sloan Digital Sky Survey V has been provided by the Alfred P. Sloan Foundation, the Heising-Simons Foundation, the National Science Foundation, and the Participating Institutions. SDSS acknowledges support and resources from the Center for High-Performance Computing at the University of Utah. The SDSS web site is www.sdss.org.
SDSS is managed by the Astrophysical Research Consortium for the Participating Institutions of the SDSS Collaboration, including the Carnegie Institution for Science, Chilean National Time Allocation Committee (CNTAC) ratified researchers, the Gotham Participation Group, Harvard University, Heidelberg University, The Johns Hopkins University, L’Ecole polytechnique fédérale de Lausanne (EPFL), Leibniz-Institut für Astrophysik Potsdam (AIP), Max-Planck-Institut für Astronomie (MPIA Heidelberg), Max-Planck-Institut für Extraterrestrische Physik (MPE), Nanjing University, National Astronomical Observatories of China (NAOC), New Mexico State University, The Ohio State University, Pennsylvania State University, Smithsonian Astrophysical Observatory, Space Telescope Science Institute (STScI), the Stellar Astrophysics Participation Group, Universidad Nacional Autónoma de México, University of Arizona, University of Colorado Boulder, University of Illinois at Urbana-Champaign, University of Toronto, University of Utah, University of Virginia, Yale University, and Yunnan University.

### v3

@@
 Abstract
-We compute parametric measurements of the Einstein-radius-enclosed total mass for 177 cluster-scale strong gravitational lenses identified by the ChicagO Optically-selected Lenses Located At the Margins of Public Surveys (COOL-LAMPS) collaboration with lens redshifts ranging from $\sim 0.2$ to $\sim 1.0$ using only two parameters: the Einstein radius and the brightest-cluster-galaxy (BCG) redshift.
-We constrain the Einstein-radius-enclosed luminosity and stellar mass by fitting parametric spectral energy distributions (SEDs) with aperture photometry from the Dark Energy Camera Legacy Survey (DECaLS) in the $grz$ Dark Energy Camera (DECam) filters.
+We compute parametric measurements of the Einstein-radius-enclosed total mass for 177 cluster-scale strong gravitational lenses identified by the ChicagO Optically-selected Lenses Located At the Margins of Public Surveys (COOL-LAMPS) collaboration with lens redshifts ranging from $0.2\lessapprox z\lessapprox 1.0$ using only two measured parameters in each lensing system: the Einstein radius, and the brightest-cluster-galaxy (BCG) redshift.
+We then constrain the Einstein-radius-enclosed luminosity and stellar mass by fitting parametric spectral energy distributions (SEDs) with aperture photometry from the Dark Energy Camera Legacy Survey (DECaLS) in the $g$, $r$, and $z$-band Dark Energy Camera (DECam) filters.
 We find that the BCG redshift, enclosed total mass, and enclosed luminosity are strongly correlated and well described by a planar relationship in 3D space.
 We also find that the enclosed total mass and stellar mass are correlated with a logarithmic slope of $0.443\pm 0.035$, and the enclosed total mass and stellar-to-total mass fraction are correlated with a logarithmic slope of $-0.563\pm 0.035$.
-The correlations described here can be used to validate strong lensing candidates in upcoming imaging surveys—such as Rubin/Legacy Survey of Space and Time (LSST)—in which an algorithmic treatment of lensing systems will be needed due to the sheer volume of data these surveys will produce.
+The correlations described here can be used to validate strong-lensing candidates in upcoming imaging surveys—such as Rubin/Legacy Survey of Space and Time (LSST)—in which an algorithmic treatment of lensing systems will be needed due to the sheer volume of data these surveys will produce.
 Galaxy clusters (584) — High-redshift galaxy clusters (2007) — Scaling relations (2031) — Spectral energy distribution (2129) — Strong gravitational lensing (1643)
@@
 Strong gravitational lensing is a rare phenomenon in the universe in which an intervening object with sufficient surface mass density bends light from a background source relative to the observer such that multiple images of the source are formed.
-For cases in which the background source is a galaxy, the lensing effect can create highly distorted and magnified images of the source galaxy in an arclike shape (e.g., Lynds & Petrosian 1986; Soucail et al.
+For cases where the background source is a galaxy, the lensing effect can create highly distorted and magnified images of the source galaxy in an arclike shape (e.g., Lynds & Petrosian 1986; Soucail et al.
 1987; Kneib & Natarajan 2011; Meneghetti et al.
@@
 2017; Shajib et al.
-2022; Zhang et al.
-2023).
+2022).
 The discovery of new gravitational lenses of this type is accelerating rapidly, particularly with the help of machine learning applied to finding galaxy-galaxy strong lenses in large optical-imaging surveys (e.g., Huang et al.
@@
 2022; Zaborowski et al.
-2022).
+2023).
 Coupled with access to a flood of new data from large surveys such as Rubin/LSST in the coming years, the population of candidate lenses both in the bulk and at the margins will only grow.
-Contemporary methodologies often rely on visual inspection in order to winnow candidate lists to an acceptable number worthy of detailed follow-up (Diehl et al., 2017); such inspection has been successful in locating new systems thus far (Rojas et al., 2023).
+Contemporary methodologies often rely on visual inspection in order to winnow candidate lists to an acceptable number worthy of detailed follow-up (e.g., Diehl et al.
+2017); such inspection has been successful in locating new systems thus far (e.g., Rojas et al.
+2023).
 However, with thousands to millions of lines of sight to choose from in an ever-growing plethora of imaging data, statistical measures and correlations in existing lensing systems may help hone candidate-lens samples by quantifying instances of strong lensing in tandem with traditional morphological identification.
@@
 2019) images, and we investigate the scaling relations between the BCG redshift, Einstein-radius-enclosed (enclosed) total mass, enclosed luminosity, enclosed stellar mass, and enclosed stellar-to-total mass fraction therein.
-The sample we use here consists of a refined subsample of all strong-lensing candidates identified by the COOL-LAMPS team, and we consider them unambiguous and robust as informed by their visual morphology—despite a lack of definite spectroscopic confirmation.
+The sample we use here consists of a refined subsample of all strong-lensing candidates identified by COOL-LAMPS, and we consider them unambiguous and robust as informed by their visual morphology—despite a lack of definite spectroscopic confirmation.
 Throughout this paper, we adopt a $\mathrm{\Lambda CDM}$ WMAP-9 cosmology (Hinshaw et al., 2013). All photometric calculations were done in the AB magnitude system.
@@
 Theory and observations both place strong constraints on observed stellar luminosity as a function of mass (e.g., Kuiper 1938; Presotto et al.
-2014; Wang & Zhong 2018), hot intracluster gas contributes significantly to the total mass in galaxy clusters (e.g., Kellogg et al.
-1972; Forman et al.
-1972; Markevitch & Vikhlinin 2007), and both galaxies and clusters are composed of a significant fraction of dark matter (e.g., Zwicky 1933; Rubin 1986; Navarro et al.
+2014; Wang & Zhong 2018), hot intracluster gas contributes significantly to the total mass in galaxy clusters (e.g., Forman et al.
+1972; Kellogg et al.
+1972; Markevitch & Vikhlinin 2007), and both galaxies and galaxy clusters are composed of a significant fraction of dark matter (e.g., Zwicky 1933; Rubin 1986; Navarro et al.
 1996; Bryan & Norman 1998; Hradecky et al.
@@
 However, robustly measuring these mass components on a per-system basis is time and resource intensive—which is unsustainable for large samples.
-Conveniently, the existence of lensing prescribes a specific amount of mass that must exist to create the observed lensing in the first place.
+Conveniently, the existence of lensing prescribes a specific amount of mass that must exist to create the observed lensing a priori.
 Remolina González et al.
-(2020) have shown that knowing only the angle between the BCG and lensed source arc (a proxy for Einstein radius; $\theta_{E}$) and the line-of-sight geometry (redshift of the lens and the source; $z_{L}$, $z_{S}$), the core mass interior to the Einstein radius of a galaxy cluster may be measured with minimal scatter on order of $\approx 10\%$.
-With these three variables, the mass can be immediately calculated and accounts for all of the projected mass density interior to the Einstein radius.
-Measuring mass in this way also offers a well-defined aperture with which to constrain photometric properties as well.
-By directly integrating over pixels inside this aperture in flux-calibrated images, the photometry can be automatically constrained and utilized by an SED modeller to obtain measurements of enclosed luminosity and stellar mass.
-The COOL-LAMPS collaboration has constructed a large dataset of strong lensing candidates using a simple human-led ranking system (e.g., Khullar et al.
-2021).
-From this, we have collated a sample of 177 galaxy clusters and groups that exhibit robust visual evidence of strong gravitational lensing in which the primary lensed arc is roughly circular with its projected center located at or very near to the BCG.
-Any systems with complex source-galaxy morphology are less likely to be well-described by the simple symmetric dark-matter profiles that inform the parametric mass estimates described in Remolina González et al.
-(2020).
-Such complex systems were intentionally excluded from the sample considered in this work.
-Figure 1: A subset of nine strong gravitational lenses used in this work.
+(2020) have shown that knowing only the angle between the BCG and lensed source arc (a proxy for Einstein radius; $\theta_{E}$) and the line-of-sight geometry (redshift of the lens and the source; $z_{L}$, $z_{S}$), the core mass interior to the Einstein radius of a strong-lensing galaxy cluster may be measured with minimal scatter on order of $\approx 10\%$.
+With these three parameters, the mass can be immediately calculated and accounts for all of the projected mass density interior to the Einstein radius.
+Measuring mass in this way also offers a well-defined aperture with which to constrain photometric properties of the lens as well.
+By directly integrating over pixels inside this aperture in flux-calibrated images, the lens-galaxy photometry can be automatically utilized by an SED modeller to obtain measurements of enclosed luminosity and stellar mass.
+The COOL-LAMPS collaboration has constructed a large dataset of strong-lensing candidates using a simple human-led ranking system (e.g., Khullar et al.
+2021; Sukay et al.
+2022; Martinez et al.
+2023; Zhang et al.
+2023; Napier et al.
+2023b; Klein et al.
+2024).
+From this, we have collated a sample of 177 galaxy clusters and groups that exhibit robust visual evidence of strong gravitational lensing in which the primary lensed source arc is roughly circular with its projected center located at or very near to the BCG.
+Any systems with complex source-galaxy morphology are less likely to be well described by the simple symmetric dark-matter profiles that inform the parametric mass estimates described in Remolina González et al.
+(2020)—such complex systems were intentionally excluded from this work.
+Figure 1: A subset of nine strong gravitational lenses analyzed in this work.
 The lenses are arranged by increasing redshift from the top to the bottom of the figure, and they are arranged by increasing Einstein radius from the left to the right of the figure.
-Red circles denote positional constraints used to derive the Einstein radius in each of the systems.
-White circles represent the Einstein aperture for each system.
-A black ‘x’ at the center of each subplot marks the exact coordinates of the BCG.
-The lens redshifts and Einstein radii for each system are also shown in each panel.
-RGB images have equal astrometric scaling and are sourced from DECaLS LS DR9 (Dey et al., 2019) using the $z$, $r$, and $g$ bands with custom color scaling respectively.
-Figure 2: Observed-frame spectral energy distributions (in maggies) for the nine systems shown in Figure 1 after SED fitting with Prospector.
-The top plot of each subplot contains the best-fit spectrum (red line), 1024 spectra drawn from their respective posterior distribution (grey lines), observed photometry (blue points), and photometric filter transmission curves for the Dark Energy Camera (DECam, Flaugher et al.
-2015) $g$, $r$, and $z$ filters from left to right respectively (in arbitrary units).
-The bottom plot of each subplot shows the residual between the best-fit spectra and the observed photometry normalized by the standard deviation of the observed photometry.
+In each subplot, red circles denote positional constraints used to derive the Einstein radius, white circles represent the Einstein aperture, and a black ‘x’ marks the exact coordinates of the BCG.
+Respective lens redshifts and Einstein radii are also shown in each subplot.
+RGB images have equal astrometric scaling and are sourced from DECaLS LS DR9 (Dey et al., 2019) using $z$, $r$, and $g$-band imaging data with custom color scaling respectively.
+Figure 2: Observed-frame spectral energy distributions (SEDs) in maggies for the nine lenses shown in Figure 1 after SED fitting with Prospector.
+The top plot of each subplot contains the best-fit SED (red line), 1024 SEDs drawn from their respective posterior distributions (grey lines), observed lens-galaxy photometry (blue points), and photometric filter transmission curves for the Dark Energy Camera (DECam, Flaugher et al.
+2015) $g$, $r$, and $z$-band filters from left to right respectively (in arbitrary units).
+The bottom plot of each subplot shows the residual between the observed photometry and the best-fit SED—normalized by the standard deviation of the observed photometry.
 2.1 Einstein Radius
-We defined the Einstein radius for each lensing system as the radius of a circle centered on the BCG (Remolina González et al., 2020) which minimizes the total angular separation between each of the three most-readily-identifiable bright “clumps” in the tangentially-lensed arcs and the perimeter of said circle.
+We defined the Einstein radius for each system as the radius of a circle centered on the BCG (Remolina González et al., 2020) which minimizes the total angular separation between each of the three most readily identifiable bright “clumps” in the tangentially lensed arcs and the perimeter of said circle.
 Henceforth, we refer to this circular aperture centered on the BCG with radius equal to the Einstein radius as the “Einstein aperture”.
-Four points (the BCG center and three points along each arc) physically constrain the Einstein aperture, and we obtained them by tagging the point of peak surface brightness for the BCG and similar peaks in surface brightness in the clumps of each tangentially-lensed arc respectively.
+Four points (the BCG center and three points along each arc) physically constrain the Einstein aperture, and we obtained them by tagging the point of peak surface brightness for the BCG and similar peaks in surface brightness in the clumps of each tangentially lensed arc respectively.
 If three distinct peaks were unable to be found along the tangential arcs, the main arc was simply traced.
-We bootstrapped the three tangential-arc positional constraints 1000 times, and we adopted the mean and standard deviation for the resulting distribution of Einstein radii as the Einstein radius and its error respectively for each system.
+We bootstrapped the three tangential-arc positional constraints 1000 times, and we adopted the mean and standard deviation from the resulting distribution of Einstein radii as the Einstein radius and its error respectively for each system.
 A descriptive visualization sampling the range of Einstein radii and lens redshifts considered in this work is shown in Figure 1.
 2.2 Lens Redshift
-BCGs are sui generis red-sequence cluster galaxies which can also be used as proxies to infer properties of the cluster-scale dark matter halo (e.g., Oegerle & Hoessel 1991; Lauer et al.
+BCGs are sui generis red-sequence cluster galaxies which can also be used as proxies to infer properties of the cluster-scale dark-matter halo (e.g., Oegerle & Hoessel 1991; Lauer et al.
 2014).
-Since we assume that the lensing systems in this work can be well-described by a symmetric dark-matter profile, we take the redshift of the BCG as interchangeable with the systematic redshift of the cluster-scale halo.
-The lens redshift converts the measured Einstein radius to a proper transverse distance via the angular diameter distance, and we query the Sloan Digital Sky Survey Data Release 15 (SDSS DR15, Aguado et al.
+Since we assume that the lensing systems in this work can be well described by a symmetric dark-matter profile, we take the redshift of the BCG as interchangeable with the systematic redshift of the cluster-scale dark-matter halo.
+The lens redshift converts the measured Einstein radius to a proper transverse distance via the angular diameter distance, and we queried the Sloan Digital Sky Survey Data Release 15 (SDSS DR15, Aguado et al.
 2019; Kollmeier et al.
 2019) in tandem with the Legacy Survey Data Release 9 (LS DR9, Dey et al.
-2019; Duncan 2022) for each system in this work to obtain BCG redshifts.
+2019; Duncan 2022) for each system in this work to obtain lens (BCG) redshifts.
 If the centroid coordinates for each BCG corresponded to a warning-free spectroscopic redshift in SDSS DR15, we adopted that redshift as the lens redshift for that system along with its associated error.
 If a given BCG lacked any corresponding value in SDSS DR15, as is the case for the higher redshift clusters in this work and/or clusters located in SDSS-non-imaged areas, we adopted the LS DR9 photometric redshift (Zhou et al., 2023) as the lens redshift and its associated error instead.
-Summarized in Table 2, we obtained spectroscopic redshifts for 41 systems, and photometric redshifts for the remaining 136.
-Table 1: Free Parameters used in SED fitting with Prospector
-Parameter Description Priors
-$\mathrm{log}(M_{tot}/M_{\odot})$ Total stellar mass formed in dex solar masses. Top Hat: [8.0, 14.0]
-$\mathrm{log}(Z/Z_{\odot})$ Stellar metallicity in dex solar metallicity. Top Hat: [-1.0, 0.2]
-$\lambda_{2}$ Diffuse dust optical depth. Top Hat: [0.0, 2.0]
-$t_{\mathrm{age}}$ Age of the cluster in Gyr. Top Hat: [$t_{z=20}$, $t_{z_{BCG}}$]
-$\tau$ SFH e-folding time in Gyr. Top Hat: [0.1, 10]
+We obtained spectroscopic redshifts for 41 systems and photometric redshifts for the remaining 136 systems.
+Table 1: Parameters used in SED Fitting with Prospector
+Free Parameter Description Priors
+Y $\mathrm{log}(M_{tot}/M_{\odot})$ Total stellar mass formed in dex solar masses. Top Hat: [8.0, 14.0].
+Y $\mathrm{log}(Z/Z_{\odot})$ Stellar metallicity in dex solar metallicity. Top Hat: [-1.0, 0.2].
+Y $\lambda_{2}$ Diffuse dust optical depth. Top Hat: [0.0, 2.0].
+Y $t_{\mathrm{age}}$ Age of the cluster in Gyr. Top Hat: [$t_{z=20}$, $t_{z_{BCG}}$].
+Y $\tau$ SFH e-folding time in Gyr. Top Hat: [0.1, 10].
+N imf_type Initial mass function type. Chabrier (Chabrier, 2003).
+N dust_type Dust attenuation curve. Calzetti (Calzetti et al., 2000).
+N sfh Star formation history model. Delayed tau (Carnall et al., 2019).
 2.3 Source Redshift
@@
 2018).
-However, obtaining these redshifts for all identified source galaxies requires prolonged additional inquiry in tension with the appeal of a robust and efficient mass estimator.
+However, obtaining these redshifts for all identified source galaxies in a given system requires prolonged additional inquiry in tension with the appeal of an efficient mass estimator.
 Remolina González et al.
 (2020) have shown that substituting a single known source-galaxy redshift with a distribution of source-galaxy redshifts introduces a statistically insignificant uncertainty into the final distribution of mass measurements when compared to the magnitude of other systematic uncertainties.
-This distribution of lensed source redshifts is measurable (Bayliss et al., 2011a, b; Tran et al., 2022), and we simply adopt a well-described Gaussian with $\mu=2$ and $\sigma=0.2$ from Bayliss et al.
+This distribution of lensed source redshifts is measurable (e.g., Bayliss et al.
+2011a, b; Tran et al.
+2022), and we simply adopt a well-described Gaussian with $\mu=2$ and $\sigma=0.2$ from Bayliss et al.
 (2011a) in keeping with the methodology of Remolina González et al.
-(2020) as the distribution of source redshifts across the entire sample.
+(2020) as the distribution of source redshifts across the entire sample in this work.
 2.4 Photometry
-[…]
+We obtained aperture photometry for SED fitting in each system by linearly summing the dereddened flux from pixels within the Einstein aperture in $g$, $r$, and $z$-band imaging data from DECaLS LS DR9.
+Since it is critical that we do not include light from the lensed source arcs as well as any intervening stars or foreground galaxies in the measured photometry, we created two masks to exclude non-lens-galaxy regions.
+The first mask removed flux from the lensed source arcs, and the second mask removed flux from interlopers as identified by a color and magnitude screening by eye.
+Accounting for masking, this linear integration was done for each of the 1000 bootstrapped Einstein apertures derived in Section 2.1, and the resulting mean and standard deviation were adopted as the $g$, $r$, and $z$-band photometry and their errors respectively for each system.
 3 Analysis
@@
 Following Remolina González et al.
-(2020), the enclosed total mass within the Einstein aperture was calculated using Equations (1) and (2), where $D(z)$ refers to the angular diameter distance at redshift $z$, via a Monte Carlo approach.
-The enclosed total mass was computed 1000 times in which each calculation used a different Einstein radius, source redshift, and lens redshift randomly drawn from a Gaussian distribution with standard deviation equal to the uncertainty on each measurement and mean equal to the value for each measurement.
+(2020), the enclosed total mass within the Einstein aperture was calculated using Equations (1) and (2)—where $D(z)$ refers to the angular diameter distance at redshift $z$—via a Monte Carlo approach.
+The enclosed total mass was computed 1000 times for each system, in which each calculation used a different Einstein radius, source redshift, and lens redshift randomly drawn from a Gaussian distribution with mean equal to the parameter value and standard deviation equal to the parameter error.
 Also described in Remolina González et al.
-(2020) is the need to apply an empirical correction described in Equation (3) based on the “completeness” of the lensed galaxy arcs in each system, where $f(\theta_{E})$ is a cubic polynomial function specified in Table 1 of Remolina González et al.
+(2020) is the need to apply an empirical correction described in Equation (3) based on the “completeness” of the lensed source arcs in each system—where $f(\theta_{E})$ is a cubic polynomial function specified in Table 1 of Remolina González et al.
 (2020).
-This is because the systematic bias and scatter of $M_{\Sigma}(<\theta_{E})$ has a dependence on the degree of symmetry for each system in which higher Einstein-radius systems have a stronger offset.
-The correction factor aims to remove this dependence as a function of Einstein radius in non-perfectly-symmetric systems.
+This is because the systematic bias and scatter of $M_{\Sigma}(<\theta_{E})$ has a dependence on the degree of symmetry for each system, in which larger-Einstein-radius systems have a stronger offset.
+The correction factor aims to remove this dependence as a function of Einstein radius in non-perfectly symmetric systems.
 $\Sigma_{\mathrm{cr}}(z_{L},z_{S})=\frac{c^{2}}{4\pi G}\frac{D_{S}(z_{S})}{D_{L}(z_{L})D_{LS}(z_{L},z_{S})}$ (1)
@@
 $\mathrm{\ }M_{\Sigma}(<\theta_{E})=\frac{M(<\theta_{E})}{f(\theta_{E})}$ (3)
-For the purposes of applying this correction in each of the 177 lensing systems in this work, we measured the azimuthal coverage ($\phi$) by determining what percentage of the Einstein aperture was traced out by the lensed galaxy arcs in a given system.
+For the purposes of applying this correction in each of the 177 lensing systems analyzed in this work, we measured the azimuthal coverage ($\phi$)—defined as the percentage of the Einstein aperture that was traced out by the lensed source arcs in a given system.
 For example, a lens with a tangential arc stretching from an azimuthal angle of 10 to 100 degrees would have $\phi=0.25$.
-This was measured by determining what azimuthal fraction of the perimeter of the Einstein aperture was subtended by the the regions masking the source arcs described in Section 2.4.
-As was shown by Remolina González et al.
-(2020), whether or not the correction factor was needed was determined by using $\phi$ as an observable proxy for the degree of symmetry for each system.
+As was done by Remolina González et al.
+(2020), $\phi$ was used as an observable proxy for the degree of symmetry in each system to determine whether or not the correction factor was needed.
 Larger-Einstein-radius systems with a predominantly low $\phi$ tend to deviate from a symmetric dark-matter halo and thus require the empirical correction.
-For spherically-symmetric systems with a large $\phi$, the measured $M_{\Sigma}(<\theta_{E})$ is taken to be fairly unbiased, and thus the correction factor is not needed.
+For spherically symmetric systems with a large $\phi$, the measured $M_{\Sigma}(<\theta_{E})$ is taken to be fairly unbiased, and thus the correction factor is not needed.
+We obtained $\phi$ by determining the fraction of the perimeter of the Einstein aperture that was subtended by the the regions masking the lensed source arcs described in Section 2.4 in each system.
 Adopting the convention used by Remolina González et al.
-2020 for each system, if $\phi<0.5$, Equation (3) was applied.
+(2020) in their analysis, if $\phi<0.5$, Equation (3) was applied.
 If $\phi\geq 0.5$, it was not applied.
-After accounting for the empirical correction, the mean and standard deviation of the resulting 1000 measurements of $M_{\Sigma}(<\theta_{E})$ were adopted as the enclosed total mass and error respectively for each system.
+Accounting for the empirical correction, the mean and standard deviation of the 1000 measurements of $M_{\Sigma}(<\theta_{E})$ were adopted as the enclosed total mass and its error respectively for each system.
 3.2 Stellar Luminosity and Mass
-We conducted SED modelling using the photometry obtained in Section 2.4 with Prospector (Johnson et al., 2021; Conroy et al., 2009; Conroy & Gunn, 2010; Foreman-Mackey et al., 2013).
-We fit a parametric SED with three assumed parameters and five free parameters.
-We assumed a Chabrier initial mass function (IMF) (Chabrier, 2003), Calzetti attenuation curve for dust (Calzetti et al., 2000), and delayed-tau star-formation history (SFH) model (Carnall et al., 2019).
-The five free parameters that were optimized are summarized in Table 1.
-We utilized emcee as implemented in Prospector with 84 walkers for a total of 6720 iterations, taking only the last 840 iterations for each of the 84 walkers (70,560 total parameter vectors) as representing the posterior distribution for each of the variables of interest in order to eliminate the significant burn-in sequence of the fit.
-After fitting, we adopted the 50th percentile of the posterior distribution as the value for each of the five parameters, and the greater difference between the 84th-50th percentile and 50th-16th percentile values as the error.
-Observed-frame spectra for the systems shown in Figure 1 after fitting are shown in Figure 2.
-To derive luminosity, we drew random sample of 1024 chains from the obtained posterior distribution and generated an SED for each.
-We integrated over the rest-frame wavelength range of 3000Å-7000Å, since this corresponds to the longest wavelength range sampled by part of at least two bands of photometry across all BCG redshifts in this work (the lower bound of 3000Å gets redshifted out of the $\mathit{g}$ and into the $\mathit{r}$ band at high redshifts).
-After integrating a given SED to obtain flux, we converted the observed flux to a luminosity in solar luminosities using the appropriate luminosity distance and adopted the mean and standard deviation from the 1024 random draws as the luminosity and its and error respectively for each system.
+We conducted parametric SED fitting using the photometry obtained in Section 2.4 with Prospector (Conroy et al., 2009; Conroy & Gunn, 2010; Johnson et al., 2021, 2023); five free parameters and three assumed parameters parameterized each SED.
+A summary of all eight parameters used in fitting can be found in Table 1.
+Utilizing emcee (Foreman-Mackey et al., 2013) as implemented in Prospector with 84 walkers for a total of 6720 iterations, only the last 840 iterations for each of the 84 walkers (70,560 total parameter vectors) were taken as representing the posterior distribution for each of the five free parameters in each system in order to eliminate the significant burn-in sequence of the fitting process.
+After fitting, we adopted the 50th percentile of the posterior distribution as the value for each free parameter, and the greater difference between the 84th-50th percentile and 50th-16th percentile values as the error.
+Observed-frame SEDs for the systems shown in Figure 1 after SED fitting are shown in Figure 2.
+To derive luminosity, we generated new SEDs with a random sample of 1024 free-parameter vectors from the posterior distributions of the best-fit SED in each system.
+We integrated over the rest-frame wavelength interval of 3000Å to 7000Å, since this corresponds to the longest wavelength interval sampled by part of at least two bands of photometry across all BCG redshifts in this work (the lower bound of 3000Å gets redshifted out of the $\mathit{g}$ and into the $\mathit{r}$ band at high redshifts).
+This range also allows us to sample the rest-frame 4000Å break while recovering flux in longer wavelengths which are significantly more luminous in cluster galaxies.
+After integrating the SEDs, we converted flux to luminosity in solar luminosities using the appropriate luminosity distance and adopted the mean and standard deviation from the 1024 random parameter vectors as the luminosity and its error respectively for each system.
 4 Results
 As Figure 3 shows, the BCG redshift, enclosed total mass, and enclosed luminosity are correlated.
-Indeed, we note that the most massive clusters in this work contain a large amount of stars and thus also have a large luminosity.
-Using emcee, we fit a plane-of-best-fit using likelihood maximization taking into account the uncertainty in variables on all three axes to describe this correlation where X = $z_{{}_{BCG}}$, Y = $\mathrm{log}_{10}[M_{\Sigma}(<\theta_{E})/M_{\odot}]$, and Z = $\mathrm{log}_{10}[L(<\theta_{E})/L_{\odot}]$, yielding:
+Indeed, we note that the most massive clusters in this work also possess a large stellar mass and thus have a large luminosity.
+Using emcee, we constrained a plane-of-best-fit using a likelihood maximization estimation taking into account the error in data on all three axes to describe this correlation—where X = $z_{{}_{BCG}}$, Y = $\mathrm{log}_{10}[M_{\Sigma}(<\theta_{E})/M_{\odot}]$, and Z = $\mathrm{log}_{10}[L(<\theta_{E})/L_{\odot}]$—yielding:
 $Z=0.137_{-0.051}^{+0.051}X+0.522_{-0.023}^{+0.022}Y+4.207_{-0.290}^{+0.300}$ (4)
 This simple and analyic relationship derived from a sample of known strong lenses, considering only quantities derived from observations, may have utility in differentiating bona fide strong lenses from non-lenses in large samples of candidate lenses where the evidence of lensing is not nearly as robust as the sample of clusters analyzed here.
-Moreover, this correlation offers a future opportunity to test simulations against observations, particularly those that aim to generate realistic galaxy populations in dense cluster environments.
-Figure 3: A 3D visualization of the BCG redshift, enclosed total mass, and enclosed luminosity for the systems in this work plotted as black circles with a plane of best fit shown in black as well.
-Uncertainties are represented as red, blue, and green lines respectively.
+Moreover, this correlation offers a future opportunity to test simulations against observations—particularly those that aim to generate realistic galaxy populations in dense cluster environments.
+Figure 3: A 3D visualization of the BCG redshift, enclosed total mass, and enclosed luminosity for the systems in this work plotted as black circles.
+A plane-of-best-fit is also shown in black as well.
+Errors in data on each axis are represented as red, blue, and green lines respectively.
 A full accounting of all measured quantities in this work can be found in Table 2.
@@
 5.1 Scaling Relations
-Besides the enclosed total mass and luminosity, we can also infer the current enclosed stellar mass at the redshift of observation for each system using the posterior distribution of each SED fit generated with Prospector.
-This was obtained by correcting the derived enclosed stellar mass in each system by the surviving mass fraction returned by Prospector, since the amount of stellar mass that we see at the time of observation is less than the total stellar mass formed over the lifetime of the galaxy (e.g., Li et al.
-2017).
-While we note that enclosed stellar mass is more complex to infer and subject to more significant systematics when compared to the more direct measures of enclosed total mass from strong lensing geometry and luminosity, it nevertheless offers another point of comparison with those measurables.
+Besides the enclosed total mass and luminosity, we can also infer the observed enclosed stellar mass at the redshift of observation for each system using the stellar-mass posterior distribution from each SED constrained with Prospector.
+We obtained this by correcting the derived enclosed stellar mass formed in each system with the surviving mass fraction returned by Prospector.
+This is necessary because the amount of stellar mass that we see at the redshift of observation is less than the total stellar mass formed over the lifetime of the galaxy (e.g., Li et al.
+2017)—the latter being what Prospector parametrically fits.
+While we note that enclosed stellar mass is more complex to infer and subject to more significant systematics when compared to the more direct measures of enclosed total mass from strong-lensing geometry and luminosity, it nevertheless offers another point of comparison with those measurables.
 5.1.1 Total Mass — Stellar Mass
-Figure 4: The total mass contained within the Einstein aperture for each system plotted against both the enclosed stellar mass and the enclosed stellar-to-total mass fraction.
-Points in panels I and III are color-coded according to the luminosity of the system in units of solar luminosities from stellar light within the Einstein aperture as measured in the rest-frame wavelength range from 3000Å to 7000Å.
-Points in panels II and IV only represent those systems in which multiple cluster members fall within the Einstein aperture.
-Linear regression lines in all four subpanels are drawn from the posterior distribution of their respective MCMC.
-In Panel I of Figure 4, we see a clearly positive correlation between the enclosed stellar mass and the enclosed total mass for each system.
-By performing a linear regression with emcee using a likelihood maximization that accounts for uncertainties in variables on both axes, we find that the slope of this correlation is $0.443\pm 0.035$ from the posterior distribution of the MCMC fit.
-We colored the points according to their luminosity to highlight how the intrinsic differences between individual systems contribute to the overall systematic scatter.
+Figure 4: The total mass enclosed within the Einstein aperture for each system on the x-axis plotted against both the enclosed stellar mass and the enclosed stellar-to-total mass fraction on the y-axes.
+Points in Panels I and III are color coded according to the luminosity of the system in units of solar luminosities from stellar light enclosed within the Einstein aperture as measured within the rest-frame wavelength interval of 3000Å to 7000Å.
+Points in Panels II and IV only represent systems in which multiple cluster members fall within the Einstein aperture.
+Linear regression lines in all four Panels are drawn from the posterior distributions of their respective MCMC.
+In Panel I of Figure 4, we see a clearly positive correlation between the enclosed total mass and the enclosed stellar mass for all systems.
+By performing a linear regression with emcee using a likelihood maximization estimation that accounts for errors in data on both axes, we find that the slope of this correlation is $0.443\pm 0.035$ from the posterior distribution of the MCMC fit.
+We colored the points according to their luminosity to highlight how the intrinsic differences between individual systems contribute to the overall intrinsic scatter.
 In Table 2 of Kravtsov et al.
-(2018), they found the stellar mass of solely the BCG and $M_{500}$ (White, 2001) to be correlated with a logarithmic slope of $0.33\pm 0.11$.
+(2018), they found $M_{500}$ (White, 2001) and the stellar mass of solely the BCG to be correlated with a logarithmic slope of $0.33\pm 0.11$.
 In addition, Kravtsov et al.
-(2018) found the stellar mass of the entire cluster and $M_{500}$ to be correlated with a logarithmic slope of $0.59\pm 0.08$.
-The fact that the slope for this work falls between the two measurements quoted by Kravtsov et al.
-(2018) is consistent with how our methodology samples a physical regime between exclusively the BCG and the entire cluster.
+(2018) found $M_{500}$ and the stellar mass of the entire cluster to be correlated with a logarithmic slope of $0.59\pm 0.08$.
+The fact that the slope for this work falls between the two slopes quoted by Kravtsov et al.
+(2018) is consistent with how our methodology samples a physical regime between exclusively the BCG and that of the entire cluster.
 Panel II of Figure 4 further explores this correlation, with only those clusters where multiple cluster members were contained within the Einstein aperture are plotted.
@@
 5.1.2 Total Mass — Stellar-to-Total Mass Fraction
-In Panel III of Figure 4, we compare the enclosed stellar-to-total mass fraction against the enclosed total mass and see a clearly negative correlation (again with the same colormapping as Panel I to emphasize intrinsic scatter).
-We find that the slope of this correlation is $-0.563\pm 0.035$.
-Andreon (2010) measured 52 clusters over a slightly larger mass range and found that the stellar mass fraction depended on halo mass with a slope of $-0.55\pm 0.08$ when measuring to a physical extent of $R_{200}$.
+In Panel III of Figure 4, we plot the enclosed stellar-to-total mass fraction against the enclosed total mass and see a clearly negative correlation (again with the same colormapping as Panel I to emphasize intrinsic scatter).
+Also using emcee, we find that the slope of this correlation is $-0.563\pm 0.035$.
 Categorizing in Panel IV identically to Panel II, we find that the multiple-galaxy systems are correlated with a slope of $-0.561\pm 0.050$.
-We again find that the multiple-galaxy systems are statistically indistinguishable from the slope for the entire sample shown in Panel III and in Andreon (2010).
-While the linear fit derived by Andreon (2010) has a slightly larger normalization, this is to be expected because the cluster masses are constrained out to $R_{200}$.
+We again find that the slope for the multiple-galaxy systems are statistically indistinguishable from the slope for the entire sample shown in Panel III.
+Andreon (2010) measured 52 clusters over a slightly larger mass range and found that the stellar mass fraction depended on halo mass with a slope of $-0.55\pm 0.08$ when measuring out to a physical extent of $R_{200}$.
+Both of the slopes in Panels III and IV are statistically similar to Andreon (2010), and while the linear fit derived by Andreon (2010) has a slightly larger normalization, this is to be expected because the cluster masses are constrained out to $R_{200}$.
 5.2 Photometric Bias and Error
-Deriving light via directly integrating over the Einstein aperture presents some challenges.
-Since we identify red-sequence cluster members by using observed color as a proxy for cluster membership, it is possible that we may have included intrinsically-bluer galaxies at a higher redshift that appear to be the same color as the cluster members.
-Conversely, we may have excluded intrinsically-bluer bona fide cluster members.
+Deriving light in Section 2.4 via directly integrating over the Einstein aperture presents some challenges.
+Since we identified red-sequence cluster members by using observed color as a proxy for cluster membership, it is possible that we may have included intrinsically bluer galaxies at redshift $z>z_{{}_{BCG}}$ that appear to be the same color as the cluster members.
+Conversely, we may have excluded intrinsically bluer bona fide cluster members.
 However, star-forming cluster members have less stellar mass and are intrinsically rarer in cluster cores (e.g., Dressler et al.
 1997, 2004), and we do not expect that these missed cluster galaxies introduce a significant loss to the measured cluster light.
-In addition, bluer galaxies at a higher redshift appear intrinsically fainter, which are also unlikely to bias our methodology of measuring cluster light.
-Interlopers such as foreground stars that appear in front of cluster galaxies are forced to be masked, but any underlying galaxy flux is typically minimal due to the limited angular size of such stars.
+In addition, bluer galaxies at redshift $z>z_{{}_{BCG}}$ appear relatively fainter, which are also unlikely to bias our methodology of measuring cluster light.
+Interlopers such as foreground stars that appear in front of cluster galaxies are forced to be masked, but any underlying lens-galaxy flux is typically minimal due to the limited angular size of such stars.
 In general, we expect that all of the unmasked flux from cluster cores at the angular scales considered in this work originates from genuine cluster members.
 Any modest background correction that could be made will also be effectively captured in the redshift dependence of the fits (e.g., Figure 3).
-While we must also consider associated galaxy populations in the immediate large-scale structure when considering galaxy clusters projected onto the sky—as such structures may also be projected onto the cluster core—those galaxies will also contribute to the lensing effect and will have their mass and light effectively captured by the methodology presented here.
-There are also extremely few examples of systems whose Einstein radius is less than $\sim 2^{\prime\prime}$.
-To some extent, this is a consequence of the chosen sample of galaxy clusters and groups, which we would generally expect to have larger Einstein radii.
-Systems whose arcs are separated from the central galaxy at small radii may be located directly within the discernable BCG light.
+While we must also consider associated galaxy populations in the immediate large-scale structure when considering galaxy clusters projected onto the sky—as such structures may also be projected onto the cluster core—those galaxies will also contribute to the lensing effect and will have their mass and light captured by the methodology presented here.
+We also note that our sample contains extremely few examples of systems which have a measured Einstein radius of $\theta_{E}\lessapprox 2^{\prime\prime}$.
+To some extent, this is a consequence of our chosen sample of cluster-scale lenses, which we would generally expect to have larger Einstein radii.
+However, we caution that there exists a large scatter between the total halo mass and Einstein radius (Fox et al., 2022).
+Systems whose arcs are separated from the central galaxy at small radii may also be located directly within the discernable BCG light.
 This would not only make them hard to visually notice, but it would also make them unsuitable for this analysis because the source light would be significantly blended together with the lensing cluster light.
-However, we caution that there is a large scatter between the total halo mass and Einstein radius (Fox et al., 2022).
-While we thus expect this work to undersample the small-Einstein-radius regime, it is not apparent that this significantly alters any of our results.
-In addition, our use of relatively shallow ground-based imaging limits the $5\sigma$ point-source magnitude depth of objects in the $grz$ bands to $\sim$ 24.7, 23.9, and 23.0 respectively in the average case of two exposures per filter (Dey et al., 2019).
-Clusters higher than redshift $\sim 1$ are barely visible even in $\mathit{z}$-band filter images in the current data and are simply a rarer type of lensing system in the universe (Li et al., 2019).
-Both the data depth and the intrinsic redshift distribution of lenses shape the redshift distribution in this work.
+While we thus expect this work to undersample the small-Einstein-radius regime, it is not apparent that this significantly alters our findings.
+In addition, our use of relatively shallow ground-based imaging limits the $5\sigma$ point-source magnitude depth of objects in the $g$, $r$, and $z$-band filters to $\approx$ 24.7, 23.9, and 23.0 respectively in the average case of two exposures per filter (Dey et al., 2019).
+Systems with a lens redshift $z\gtrapprox 1$ are barely visible even in $\mathit{z}$-band LS DR9 imaging data, and they are also simply a rarer type of lensing system in the universe (Li et al., 2019).
+Both the data depth and intrinsic redshift distribution of lenses shape the redshift distribution in this work.
 5.3 Alternative Mass Measurement
-In constraining mass, we also considered a second method for measuring the cluster-centric mass elucidated in Remolina González et al.
-(2021a) using the parametric lens modelling software LENSTOOL (Jullo et al., 2007) to measure the enclosed mass as opposed to the simple evaluation of Equations (1), (2), and (3).
-For the first 35 clusters that were analyzed, we generated single-halo lens models by using the three coordinates corresponding to the locations of the brightest clumps in each lensed arc as positional constraints with a single pseudo-isothermal elliptical mass distribution (PIEMD, Kassiola & Kovner 1993) locked to the center of the BCG.
-Mass estimates were then obtained from the resulting best-fit models according to the methodology described in Remolina González et al.
+In constraining mass, we also considered a second method for measuring cluster-centric mass elucidated in Remolina González et al.
+(2021a) using the parametric lens-modelling software LENSTOOL (Jullo et al., 2007) to measure the enclosed total mass—as opposed to the simple evaluation of Equations (1), (2), and (3).
+For the first 35 clusters that were analyzed in this work, we generated single-halo lens models by taking the three coordinates in each lensed arc used to derive the Einstein radius in Section 2.1 as a multiple-image family with a single pseudo-isothermal elliptical mass distribution (PIEMD, Kassiola & Kovner 1993) locked to the center of the BCG.
+Mass estimates were then obtained from the resulting best-fit lens models according to the methodology described in Remolina González et al.
 (2021a).
-Figure 5: A comparison of the single-halo LENSTOOL generated mass estimates ($M_{SHM}$) and the parametrically generated mass estimates ($M_{\Sigma}$) for the first 35 clusters analyzed in this work.
-The black line represents an ideal 1-1 correspondence, and the blue lines represent the uncertainty on the real correspondence.
+Figure 5: A comparison of the single-halo LENSTOOL-generated mass measurements ($M_{SHM}$) and the parametrically generated mass measurements ($M_{\Sigma}$) for the first 35 clusters analyzed in this work.
+The black line represents an ideal 1-1 correspondence, and the blue lines represent the uncertainty on the measured correspondence.
 As expected from Remolina González et al.
-(2021b), the two different methodologies show similar results, and the two mass measurements for the 35 clusters studied here are correlated with a slope of $0.960\pm 0.036$ and intercept of $0.492\pm 0.476$ as shown in Figure 5.
-We abandoned the LENSTOOL approach throughout the rest of the work, i.e.
-we find no evidence for a significant difference between these two methods.
+(2021b), these differing methodologies show similar results.
+The two mass measurements for the 35 clusters studied here are correlated with a slope of $0.960\pm 0.036$ and intercept of $0.492\pm 0.476$ dex as shown in Figure 5.
+In other words, we find no reason to suggest a preference between one method over the other, and we abandoned the LENSTOOL approach in favor of the much faster parametric approach throughout the rest of the work.
 In theory, lens modelling has the potential to be the most accurate mass-constraining methodology.
 However, this requires high-spatial-resolution imaging data and ample time.
-The process of iteratively refining just one lens model may take upwards of an hour for even a single relatively-simple system, which is untenable when the number of candidate systems reaches into the thousands.
-Moreover, the accuracy of measuring mass with lens modelling is sensitive to the choice of defining multiple image families along the lensed arc in a way that simple circle fitting is not.
-These essential positional constraints are often unresolved with even the highest-quality ground-based data.
+The process of iteratively refining just one lens model may take upwards of an hour for even a single relatively simple system—which is untenable when the number of candidate systems reaches into the thousands.
+Moreover, the accuracy of measuring mass with lens modelling is sensitive to the choice of defining multiple-image families along the lensed source arcs in a way that simple circle fitting is not.
+These essential positional constraints are often unresolved even with the highest-quality ground-based imaging data.
 Conversely, the positional constraints that inform the Einstein radius ultimately used by Equations (1), (2), and (3) need not be extremely accurate with respect to exactly constraining image family positions.
-As long as the positional constraints are reasonably placed within the arcs, something which is resolved by ground-based data, it is accurate to within the systematic uncertainty.
+As long as the positional constraints are reasonably placed within the arcs, something which is resolved by ground-based imaging data, it is accurate to within the systematic uncertainty.
+For more information on lens-modelling software, related techniques, and applications of strong lensing, we refer the reader to the following non-exhaustive list of papers for further review: Elíasdóttir et al.
+(2007), Oguri (2010), Lefor et al.
+(2013), Meneghetti et al.
+(2017), Birrer & Amara (2018), Schäfer et al.
+(2020), Sharon et al.
+(2022), and Napier et al.
+(2023a).
 6 Summary and Conclusions
@@
 By directly obtaining the Einstein radius and BCG redshift for each system, we compute the Einstein-radius-enclosed total mass parametrically.
-We use the enclosed $grz$ photometry and Prospector to parameterize the enclosed luminosity and stellar mass, and we obtain a planar relationship between the BCG redshift, enclosed total mass, and enclosed luminosity of the form $Z=0.137_{-0.051}^{+0.051}X+0.522_{-0.023}^{+0.022}Y+4.207_{-0.290}^{+0.300}$ where X = $z_{{}_{BCG}}$, Y = $\mathrm{log}_{10}[M_{\Sigma}(<\theta_{E})/M_{\odot}]$, and Z = $\mathrm{log}_{10}[L(<\theta_{E})/L_{\odot}]$.
-We find that the enclosed total mass and stellar mass for the entire sample of strong lensing systems considered in this work are correlated with a logarithmic slope, normalization, and scatter of $0.443\pm 0.035$, $5.958\pm 0.454$, and $0.173\pm 0.013$ respectively.
-We also find that the enclosed total mass and stellar-to-total mass fraction are correlated with a logarithmic slope, normalization, and scatter of $-0.563\pm 0.035$, $6.034\pm 0.457$, and $0.161\pm 0.013$ respectively.
+We use the enclosed $g$, $r$, and $z$-band photometry with Prospector to parameterize the enclosed luminosity and stellar mass, and we obtain a planar relationship between the BCG redshift, enclosed total mass, and enclosed luminosity of the form $Z=0.137_{-0.051}^{+0.051}X+0.522_{-0.023}^{+0.022}Y+4.207_{-0.290}^{+0.300}$ where X = $z_{{}_{BCG}}$, Y = $\mathrm{log}_{10}[M_{\Sigma}(<\theta_{E})/M_{\odot}]$, and Z = $\mathrm{log}_{10}[L(<\theta_{E})/L_{\odot}]$.
+We find that the enclosed total mass and stellar mass for the entire sample of strong-lensing systems considered in this work are correlated with a logarithmic slope, normalization, and intrinsic stellar-mass scatter of $0.443\pm 0.035$, $5.958\pm 0.454$ dex, and $0.173\pm 0.013$ dex respectively.
+We also find that the enclosed total mass and stellar-to-total mass fraction are correlated with a logarithmic slope, normalization, and intrinsic stellar-to-total-mass-fraction scatter of $-0.563\pm 0.035$, $6.034\pm 0.457$ dex, and $0.161\pm 0.013$ dex respectively.
 Enormous volumes of imaging data from new space-based and ground-based surveys at deeper and sharper photometric limits are on the horizon.
 Probing these upcoming datasets with both visual and machine-learning driven searches for new strong lenses at a variety of scales is inevitable.
-The correlations shown in this work should have utility in filtering rankings of possible strong lensing systems as well as offering another measure of cluster cores that can be compared to cosmological simulations.
+The correlations shown in this work should have utility in filtering rankings of possible strong-lensing systems as well as offering another measure of cluster cores that can be compared to cosmological simulations.
 Acknowledgments

### v4

@@
-COOL-LAMPS. VII. Quantifying Strong-lens Scaling Relations with 177 Cluster-scale Gravitational Lenses in DECaLS
+COOL-LAMPS. VII. Quantifying Strong-lens Scaling Relations with 177 Cluster-scale Strong Gravitational Lenses in DECaLS
 Abstract
-We compute parametric measurements of the Einstein-radius-enclosed total mass for 177 cluster-scale strong gravitational lenses identified by the ChicagO Optically-selected Lenses Located At the Margins of Public Surveys (COOL-LAMPS) collaboration with lens redshifts ranging from $0.2\lessapprox z\lessapprox 1.0$ using only two measured parameters in each lensing system: the Einstein radius, and the brightest-cluster-galaxy (BCG) redshift.
-We then constrain the Einstein-radius-enclosed luminosity and stellar mass by fitting parametric spectral energy distributions (SEDs) with aperture photometry from the Dark Energy Camera Legacy Survey (DECaLS) in the $g$, $r$, and $z$-band Dark Energy Camera (DECam) filters.
+We estimate the Einstein-radius-enclosed total mass for 177 cluster-scale strong gravitational lenses identified by the ChicagO Optically selected Lenses Located At the Margins of Public Surveys (COOL-LAMPS) collaboration with lens redshifts ranging from $0.2\lessapprox z\lessapprox 1.0$ using the brightest-cluster-galaxy (BCG) redshift and an observable proxy for the Einstein radius.
+We constrain the Einstein-radius-enclosed luminosity and stellar mass by fitting parametric spectral energy distributions to aperture photometry from the Dark Energy Camera Legacy Survey in the $g$-, $r$-, and $z$-band Dark Energy Camera filters.
 We find that the BCG redshift, enclosed total mass, and enclosed luminosity are strongly correlated and well described by a planar relationship in 3D space.
-We also find that the enclosed total mass and stellar mass are correlated with a logarithmic slope of $0.443\pm 0.035$, and the enclosed total mass and stellar-to-total mass fraction are correlated with a logarithmic slope of $-0.563\pm 0.035$.
-The correlations described here can be used to validate strong-lensing candidates in upcoming imaging surveys—such as Rubin/Legacy Survey of Space and Time (LSST)—in which an algorithmic treatment of lensing systems will be needed due to the sheer volume of data these surveys will produce.
-Galaxy clusters (584) — High-redshift galaxy clusters (2007) — Scaling relations (2031) — Spectral energy distribution (2129) — Strong gravitational lensing (1643)
+We find that the enclosed total mass and stellar mass are correlated with a logarithmic slope of $0.500^{+0.029}_{-0.031}$, and the enclosed total mass and stellar-to-total mass fraction are correlated with a logarithmic slope of $-0.495^{+0.032}_{-0.033}$.
+In tandem with the small radii within which these slopes are constrained, this may suggest invariance in baryon conversion efficiency and feedback strength as a function of cluster-centric radii in galaxy clusters.
+Additionally, the correlations described here should have utility in ranking strong-lensing candidates in upcoming imaging surveys—such as Rubin/Legacy Survey of Space and Time—in which an algorithmic treatment of strong lenses will be needed due to the sheer volume of data these surveys will produce.
+Galaxy clusters (584) ; High-redshift galaxy clusters (2007) ; Scaling relations (2031) ; Spectral energy distribution (2129) ; Strong gravitational lensing (1643)
+journal: ApJfacilities: Sloan (BOSS spectrograph), Blanco (DECam wide-field camera)software: astropy (Astropy Collaboration et al., 2013, 2018, 2022), astro-prospector (Johnson et al., 2021), emcee (Foreman-Mackey et al., 2013), FSPS (Conroy et al., 2009; Conroy & Gunn, 2010), Jupyter Notebook (Kluyver et al., 2016), LENSTOOL (Jullo et al., 2007), matplotlib (Hunter, 2007), numpy (Harris et al., 2020), pandas (McKinney, 2010), python-FSPS (Johnson et al., 2023), SAOImageDS9 (Joye & Mandel, 2003), scipy (Virtanen et al., 2020).
 […]
+[…]
+2025 January 27
 1 Introduction
-Strong gravitational lensing is a rare phenomenon in the universe in which an intervening object with sufficient surface mass density bends light from a background source relative to the observer such that multiple images of the source are formed.
-For cases where the background source is a galaxy, the lensing effect can create highly distorted and magnified images of the source galaxy in an arclike shape (e.g., Lynds & Petrosian 1986; Soucail et al.
+Figure 1: A normalized histogram of the fitted Einstein radii in log space for this work shown in blue.
+Overplotted in black is a normalized Gaussian with one-sided Gaussian roll-off fit to the cumulative distribution function (CDF) of these radii in dex (see Section 2.1 for parameters describing this fit).
+Also plotted in red is a Gaussian probability distribution function with $\mu=0.73$ and $\sigma=0.316$ from Zitrin et al.
+(2012) representing the simulated Einstein radius distribution of 10,000 Sloan Digital Sky Survey (SDSS) clusters.
+Strong gravitational lensing is a rarely observed phenomenon in the Universe in which an intervening object with sufficient surface-mass density bends light from a background source relative to the observer such that multiple images of the source are formed.
+For cases where the background source is a galaxy, the lensing effect can create highly distorted and magnified images of the source galaxy in an arc-like shape (e.g., Lynds & Petrosian 1986; Soucail et al.
 1987; Kneib & Natarajan 2011; Meneghetti et al.
@@
 2017; Shajib et al.
-2022).
-The discovery of new gravitational lenses of this type is accelerating rapidly, particularly with the help of machine learning applied to finding galaxy-galaxy strong lenses in large optical-imaging surveys (e.g., Huang et al.
+2024).
+The discovery of new strong lenses is accelerating rapidly, particularly with the help of machine learning applied to large optical-imaging surveys (e.g., Huang et al.
 2021; Rojas et al.
@@
 2023).
-Coupled with access to a flood of new data from large surveys such as Rubin/LSST in the coming years, the population of candidate lenses both in the bulk and at the margins will only grow.
-Contemporary methodologies often rely on visual inspection in order to winnow candidate lists to an acceptable number worthy of detailed follow-up (e.g., Diehl et al.
+Coupled with access to a flood of new data from large surveys such as Rubin/Legacy Survey of Space and Time (LSST) in the coming years, the population of candidate lenses both in the bulk and at the margins will only grow.
+Any flux-limited survey will have tentative arcs; i.e.
+at the flux limit.
+Regardless of the lensing configuration, numerous faint features that are suggestive of strong lensing, but whose reality and interpretation as strong lensing, will exist and require verification.
+Contemporary methodologies of identifying strong lenses often rely on visual inspection in order to winnow candidate lists to an acceptable number worthy of detailed follow-up (e.g., Diehl et al.
 2017); such inspection has been successful in locating new systems thus far (e.g., Rojas et al.
 2023).
-However, with thousands to millions of lines of sight to choose from in an ever-growing plethora of imaging data, statistical measures and correlations in existing lensing systems may help hone candidate-lens samples by quantifying instances of strong lensing in tandem with traditional morphological identification.
-This would produce a purer sample and help streamline the ability to identify targets.
-In this paper, we analyze a sample of 177 strong gravitational lenses identified via visual inspection by the COOL-LAMPS collaboration in DECaLS Legacy Survey Data Release 8 (LS DR8, Dey et al.
-2019) images, and we investigate the scaling relations between the BCG redshift, Einstein-radius-enclosed (enclosed) total mass, enclosed luminosity, enclosed stellar mass, and enclosed stellar-to-total mass fraction therein.
-The sample we use here consists of a refined subsample of all strong-lensing candidates identified by COOL-LAMPS, and we consider them unambiguous and robust as informed by their visual morphology—despite a lack of definite spectroscopic confirmation.
-Throughout this paper, we adopt a $\mathrm{\Lambda CDM}$ WMAP-9 cosmology (Hinshaw et al., 2013). All photometric calculations were done in the AB magnitude system.
+In addition, automated strong-lens modeling has emerged as a promising avenue toward blindly identifying multiple-image families in strong lenses (e.g., Stapelberg et al.
+2019; Carrasco et al.
+2020; Zalesky & Ebeling 2020), though such methodologies typically rely on a light-traces-mass model to feed to a preliminary lens model—of which the latter can be computationally expensive.
+With thousands to millions of lines of sight to choose from in an ever-growing plethora of imaging data, analyzing statistical measures and correlations in existing strong lenses may help hone candidate-lens samples by quantifying instances of cluster-scale strong lensing in tandem with geometric identification.
+Figure 2: A subset of nine strong gravitational lenses analyzed in this work.
+The lenses are arranged by increasing redshift from the top to the bottom of the Figure, and they are arranged by increasing Einstein radius from the left to the right of the Figure.
+Respective lens redshifts and Einstein radii are also shown in each subplot.
+RGB images have equal astrometric scaling and are sourced from DECaLS LS DR9 (Dey et al., 2019) using $z$-, $r$-, and $g$-band imaging data with custom color scaling, respectively.
+This Figure is available as an animation.
+The animation fades in and out red circles, which denote positional constraints used to derive the Einstein radius, white circles representing the Einstein aperture, and a black ‘x’ marking the exact coordinates of each BCG.
+(An animation of this figure is available in the online article.)
+This work aims to quantify scaling relations in existing cluster-scale strong lenses such that ambiguous systems could be analyzed with a similar methodology to determine if a given system lies on or off the scaling relationships we derive here.
+Coupled with existing algorithms that can identify preferential lines of sight for cluster-scale strong lensing (e.g., Wong et al.
+2013) and an arc-finding algorithm (e.g., Seidel & Bartelmann 2007), these calculations could be undertaken automatically for arbitrary imaging data.
+This analysis would help streamline the ability to identify targets for follow-up observations by ranking candidate lenses to produce a purer sample.
+We analyze a sample of 177 cluster-scale strong gravitational lenses identified via visual inspection by the COOL-LAMPS collaboration in DECaLS Legacy Survey Data Release 8 (LS DR8, Dey et al.
+2019) images, and we investigate the scaling relations between the brightest-cluster-galaxy (BCG) redshift, Einstein-radius-enclosed (enclosed) total mass, enclosed luminosity, enclosed stellar mass, and enclosed stellar-to-total mass fraction therein.
+The sample we use here consists of a refined subsample of all strong-lensing candidates identified by COOL-LAMPS that we consider robust—despite a lack of definite spectroscopic confirmation in many cases.
+All of the lenses used here have the visual geometry of a giant arc with clear curvature that is bright enough to be unambiguously visible in the survey data.
+Figure 3: Observed-frame spectral energy distributions (SEDs) in maggies for the nine lenses shown in Figure 2 after SED fitting with Prospector.
+The top plot of each subplot contains the best-fit SED (red line), 1024 SEDs drawn from their respective posterior distributions (gray lines), observed cluster photometry (blue points), and photometric filter transmission curves for the Dark Energy Camera (DECam; Flaugher et al.
+2015) $g$-, $r$-, and $z$-band filters from left to right, respectively (in arbitrary units).
+The bottom plot of each subplot shows the residual between the observed photometry and the best-fit SED—normalized by the standard deviation of the observed photometry.
+Throughout this paper, we adopt a flat cold dark matter WMAP-9 cosmology (Hinshaw et al., 2013). All photometric calculations were done in the AB magnitude system.
 2 Methodology
@@
 Theory and observations both place strong constraints on observed stellar luminosity as a function of mass (e.g., Kuiper 1938; Presotto et al.
-2014; Wang & Zhong 2018), hot intracluster gas contributes significantly to the total mass in galaxy clusters (e.g., Forman et al.
+2014; Wang & Zhong 2018), hot intracluster gas and diffuse intracluster light contribute to the total mass in galaxy clusters (e.g., Forman et al.
 1972; Kellogg et al.
-1972; Markevitch & Vikhlinin 2007), and both galaxies and galaxy clusters are composed of a significant fraction of dark matter (e.g., Zwicky 1933; Rubin 1986; Navarro et al.
+1972; Gonzalez et al.
+2005; Markevitch & Vikhlinin 2007), and both galaxies and galaxy clusters are composed of a significant fraction of dark matter (e.g., Zwicky 1933; Rubin 1986; Navarro et al.
 1996; Bryan & Norman 1998; Hradecky et al.
@@
 Remolina González et al.
-(2020) have shown that knowing only the angle between the BCG and lensed source arc (a proxy for Einstein radius; $\theta_{E}$) and the line-of-sight geometry (redshift of the lens and the source; $z_{L}$, $z_{S}$), the core mass interior to the Einstein radius of a strong-lensing galaxy cluster may be measured with minimal scatter on order of $\approx 10\%$.
-With these three parameters, the mass can be immediately calculated and accounts for all of the projected mass density interior to the Einstein radius.
-Measuring mass in this way also offers a well-defined aperture with which to constrain photometric properties of the lens as well.
-By directly integrating over pixels inside this aperture in flux-calibrated images, the lens-galaxy photometry can be automatically utilized by an SED modeller to obtain measurements of enclosed luminosity and stellar mass.
-The COOL-LAMPS collaboration has constructed a large dataset of strong-lensing candidates using a simple human-led ranking system (e.g., Khullar et al.
+(2020) have shown that the core mass interior to the Einstein radius of a strong-lensing galaxy cluster may be computed with an uncertainty of $\approx 10\%$ via the simple evaluation of a function parameterized by the angle between the BCG and lensed source arc (a proxy for Einstein radius; $\theta_{E}$) and the line-of-sight geometry (redshift of the lens and the source; $z_{L}$, $z_{S}$).
+With these three parameters, the mass can be immediately calculated and accounts for all of the projected mass density interior to the Einstein radius for a strong lens whose dark matter halo can be described by a circularly symmetric dark matter profile.
+Measuring mass in this way also offers a well-defined aperture with which to constrain photometric properties of the strong lens as well.
+By directly integrating over pixels that correspond to cluster members inside this aperture in flux-calibrated images, the photometry of the strong lens can be automatically utilized by a spectral energy distribution (SED) modeler to obtain measurements of enclosed luminosity and stellar mass.
+The COOL-LAMPS collaboration has constructed and analyzed a large dataset of strong-lensing candidates using a simple human-led ranking system (e.g., Khullar et al.
 2021; Sukay et al.
@@
 2023b; Klein et al.
+2024; Cloonan et al.
 2024).
-From this, we have collated a sample of 177 galaxy clusters and groups that exhibit robust visual evidence of strong gravitational lensing in which the primary lensed source arc is roughly circular with its projected center located at or very near to the BCG.
-Any systems with complex source-galaxy morphology are less likely to be well described by the simple symmetric dark-matter profiles that inform the parametric mass estimates described in Remolina González et al.
-(2020)—such complex systems were intentionally excluded from this work.
-Figure 1: A subset of nine strong gravitational lenses analyzed in this work.
-The lenses are arranged by increasing redshift from the top to the bottom of the figure, and they are arranged by increasing Einstein radius from the left to the right of the figure.
-In each subplot, red circles denote positional constraints used to derive the Einstein radius, white circles represent the Einstein aperture, and a black ‘x’ marks the exact coordinates of the BCG.
-Respective lens redshifts and Einstein radii are also shown in each subplot.
-RGB images have equal astrometric scaling and are sourced from DECaLS LS DR9 (Dey et al., 2019) using $z$, $r$, and $g$-band imaging data with custom color scaling respectively.
-Figure 2: Observed-frame spectral energy distributions (SEDs) in maggies for the nine lenses shown in Figure 1 after SED fitting with Prospector.
-The top plot of each subplot contains the best-fit SED (red line), 1024 SEDs drawn from their respective posterior distributions (grey lines), observed lens-galaxy photometry (blue points), and photometric filter transmission curves for the Dark Energy Camera (DECam, Flaugher et al.
-2015) $g$, $r$, and $z$-band filters from left to right respectively (in arbitrary units).
-The bottom plot of each subplot shows the residual between the observed photometry and the best-fit SED—normalized by the standard deviation of the observed photometry.
+From this, we have collated a sample of 177 galaxy clusters that exhibit robust visual evidence of strong gravitational lensing in which the primary lensed source arc is roughly circular with its projected center located at or very near to the BCG.
+Any systems with complex source-galaxy geometry in the image plane imply that the lens possesses a complex dark matter distribution that is less likely to be well described a circularly symmetric dark matter profile.
+Since the assumption of a circularly symmetric dark matter profile underpins the functions used to estimate the enclosed mass described in Remolina González et al.
+(2020), such complex systems were intentionally excluded from this work.
 2.1 Einstein Radius
-We defined the Einstein radius for each system as the radius of a circle centered on the BCG (Remolina González et al., 2020) which minimizes the total angular separation between each of the three most readily identifiable bright “clumps” in the tangentially lensed arcs and the perimeter of said circle.
+We defined the Einstein radius for each system as the radius of a circle centered on the BCG (Remolina González et al., 2020) that minimizes the total angular separation between each of the three most readily identifiable bright “clumps” in the tangentially lensed arcs and the perimeter of said circle.
 Henceforth, we refer to this circular aperture centered on the BCG with radius equal to the Einstein radius as the “Einstein aperture”.
-Four points (the BCG center and three points along each arc) physically constrain the Einstein aperture, and we obtained them by tagging the point of peak surface brightness for the BCG and similar peaks in surface brightness in the clumps of each tangentially lensed arc respectively.
+Four points (the BCG center and three points along each arc) physically constrain the Einstein aperture, and we obtained them by tagging the point of peak surface brightness for the BCG and similar peaks in surface brightness in the clumps of each tangentially lensed arc, respectively.
 If three distinct peaks were unable to be found along the tangential arcs, the main arc was simply traced.
-We bootstrapped the three tangential-arc positional constraints 1000 times, and we adopted the mean and standard deviation from the resulting distribution of Einstein radii as the Einstein radius and its error respectively for each system.
-A descriptive visualization sampling the range of Einstein radii and lens redshifts considered in this work is shown in Figure 1.
+Each tangential-arc point was assigned a Gaussian uncertainty in both spatial axes with $\mu$ equal to its location and $\sigma=0.5^{\prime\prime}$ which roughly corresponds to the width of an average arc.
+We bootstrapped the three tangential-arc points 1000 times, with each sampling drawn from the 2D Gaussian distribution for each point.
+We adopted the mean and standard deviation from the resulting distribution of Einstein radii as the Einstein radius and its error, respectively, for each system.
+By fitting to the cumulative distribution function (CDF) of the Einstein radii for all 177 systems in this work, we find that they are distributed log normally with $\mu=0.71\pm 0.07$ and $\sigma=0.31\pm 0.04$.
+We also parameterize an incompleteness roll-off to low values with a one-sided Gaussian centered at $0.49\pm 0.16$ with $\sigma=0.14\pm 0.07$ which indicates that our sample is incomplete starting at $3.09^{\prime\prime}\,{}^{+1.37}_{-0.99}$ and 50% incomplete by $2.11^{\prime\prime}\,{}^{+1.08}_{-0.74}$.
+Physical motivations for this incompleteness are further discussed in Section 5.2.
+Zitrin et al.
+(2012) simulated the “universal” distribution of Einstein radii for a sample of 10,000 Sloan Digital Sky Survey (SDSS) clusters and found that this distribution is log normal with $\mu=0.73^{+0.02}_{-0.03}$ and $\sigma=0.316^{+0.004}_{-0.002}$.
+As shown in Figure 1, we find good agreement between the distribution of Einstein radii in this work and that of Zitrin et al.
+(2012)—indicating that our methodology is a good proxy for the true underlying Einstein radius for our sample of galaxy clusters.
 2.2 Lens Redshift
-BCGs are sui generis red-sequence cluster galaxies which can also be used as proxies to infer properties of the cluster-scale dark-matter halo (e.g., Oegerle & Hoessel 1991; Lauer et al.
+BCGs are sui generis red-sequence cluster galaxies that can also be used as proxies to infer properties of the cluster-scale dark matter halo (e.g., Oegerle & Hoessel 1991; Lauer et al.
 2014).
-Since we assume that the lensing systems in this work can be well described by a symmetric dark-matter profile, we take the redshift of the BCG as interchangeable with the systematic redshift of the cluster-scale dark-matter halo.
-The lens redshift converts the measured Einstein radius to a proper transverse distance via the angular diameter distance, and we queried the Sloan Digital Sky Survey Data Release 15 (SDSS DR15, Aguado et al.
+Since we assume that the strong lenses in this work can be well described by a symmetric dark matter profile, we take the redshift of the BCG as interchangeable with the systematic redshift of the cluster-scale dark matter halo.
+The lens redshift converts the measured Einstein radius to a proper transverse distance via the angular diameter distance, and we queried the SDSS Data Release 15 (DR15, Aguado et al.
 2019; Kollmeier et al.
@@
 If the centroid coordinates for each BCG corresponded to a warning-free spectroscopic redshift in SDSS DR15, we adopted that redshift as the lens redshift for that system along with its associated error.
-If a given BCG lacked any corresponding value in SDSS DR15, as is the case for the higher redshift clusters in this work and/or clusters located in SDSS-non-imaged areas, we adopted the LS DR9 photometric redshift (Zhou et al., 2023) as the lens redshift and its associated error instead.
+If a given BCG lacked any corresponding value in SDSS DR15, as is the case for the higher-redshift clusters in this work and/or clusters located in SDSS-non-imaged areas, we adopted the LS DR9 photometric redshift (Zhou et al., 2023) as the lens redshift and its associated error instead.
 We obtained spectroscopic redshifts for 41 systems and photometric redshifts for the remaining 136 systems.
+A descriptive visualization sampling the range of Einstein radii and lens redshifts considered in this work is shown in Figure 2.
 Table 1: Parameters Used in SED Fitting with Prospector
@@
 Y $\lambda_{2}$ Diffuse dust optical depth. Top Hat: [0.0, 2.0].
-Y $t_{\mathrm{age}}$ Age of the cluster in Gyr. Top Hat: [$t_{z=20}$, $t_{z_{BCG}}$].
-Y $\tau$ SFH e-folding time in Gyr. Top Hat: [0.1, 10].
+Y $t_{\mathrm{age}}$ Age of the cluster in gigayears. Top Hat: [$t_{z=20}$, $t_{z_{BCG}}$].
+Y $\tau$ SFH e-folding time in gigayears. Top Hat: [0.1, 10].
 N imf_type Initial mass function type. Chabrier (Chabrier, 2003).
@@
 Remolina González et al.
-(2020) have shown that substituting a single known source-galaxy redshift with a distribution of source-galaxy redshifts introduces a statistically insignificant uncertainty into the final distribution of mass measurements when compared to the magnitude of other systematic uncertainties.
+(2020) showed that substituting a single known source-galaxy redshift with a distribution of source-galaxy redshifts introduces a statistically insignificant uncertainty into the final distribution of mass measurements when compared to the magnitude of other systematic uncertainties.
 This distribution of lensed source redshifts is measurable (e.g., Bayliss et al.
@@
 (2020) as the distribution of source redshifts across the entire sample in this work.
+We need not consider the redshift distribution of sources other than giant arcs, since all of the source galaxies we analyze in this work are geometrically consistent with the giant arcs analyzed in Bayliss et al.
+(2011a).
+While we note that a later paper by Bayliss (2012) favors a slightly smaller error with the same mean, using this distribution would only marginally impact the uncertainties on the calculated total masses.
+We therefore use the distribution in Bayliss et al.
+(2011a) to maintain continuity with Remolina González et al.
+(2020).
 2.4 Photometry
-We obtained aperture photometry for SED fitting in each system by linearly summing the dereddened flux from pixels within the Einstein aperture in $g$, $r$, and $z$-band imaging data from DECaLS LS DR9.
-Since it is critical that we do not include light from the lensed source arcs as well as any intervening stars or foreground galaxies in the measured photometry, we created two masks to exclude non-lens-galaxy regions.
+We obtained aperture photometry for SED fitting in each system by linearly summing the dereddened flux from pixels within the Einstein aperture in $g$-, $r$-, and $z$-band imaging data from DECaLS LS DR9.
+Since it is critical that we do not include light from the lensed source arcs as well as any interloping stars or galaxies that are not a member of the lensing cluster in the measured lens photometry, we created two masks to exclude non-cluster-member regions.
 The first mask removed flux from the lensed source arcs, and the second mask removed flux from interlopers as identified by a color and magnitude screening by eye.
-Accounting for masking, this linear integration was done for each of the 1000 bootstrapped Einstein apertures derived in Section 2.1, and the resulting mean and standard deviation were adopted as the $g$, $r$, and $z$-band photometry and their errors respectively for each system.
+Accounting for masking, this linear integration was done for each of the 1000 bootstrapped Einstein apertures derived in Section 2.1.
+The resulting mean was adopted as the $g$-, $r$-, and $z$-band photometry for each system—with the standard deviation quantifying the geometric error.
+In quantifying the photometric error per-band, we first fit a sigma-clipped Gaussian to a histogram of the pixel values characterizing the background in each band for each system.
+The standard deviation of this Gaussian multiplied by $\sqrt{N}$, where $N$ is the number of pixels interior to the Einstein aperture, quantifies the statistical uncertainty in the photometric error for each band in each system.
+Next, the standard deviations for each Gaussian in a given photometric band across all systems were plotted as a histogram and fitted with another Gaussian.
+The standard deviation of this second Gaussian multiplied by $N$ characterizes the systematic uncertainty of the sky subtraction for each band across all systems.
+Taken together, the quadrature sum of the geometric Einstein aperture error, statistical photometric error, and systematic sky-subtraction error were adopted as the $g$-, $r$-, and $z$-band error for each system.
 3 Analysis
@@
 (2020).
-This is because the systematic bias and scatter of $M_{\Sigma}(<\theta_{E})$ has a dependence on the degree of symmetry for each system, in which larger-Einstein-radius systems have a stronger offset.
+This is because the systematic bias and scatter of $M_{\Sigma}(<\theta_{E})$ has a dependence on the degree of symmetry for each system—in which larger-Einstein-radius systems have a stronger offset.
 The correction factor aims to remove this dependence as a function of Einstein radius in non-perfectly symmetric systems.
@@
 $\mathrm{\ }M_{\Sigma}(<\theta_{E})=\frac{M(<\theta_{E})}{f(\theta_{E})}$ (3)
-For the purposes of applying this correction in each of the 177 lensing systems analyzed in this work, we measured the azimuthal coverage ($\phi$)—defined as the percentage of the Einstein aperture that was traced out by the lensed source arcs in a given system.
-For example, a lens with a tangential arc stretching from an azimuthal angle of 10 to 100 degrees would have $\phi=0.25$.
+For the purposes of applying this correction in each of the 177 strong lenses analyzed in this work, we measured the azimuthal coverage ($\phi$)—defined as the percentage of the Einstein aperture that was traced out by the lensed source arcs in a given system.
+For example, a lens with a tangential arc stretching from an azimuthal angle of $10^{\circ}$ to $100^{\circ}$ would have $\phi=0.25$.
 As was done by Remolina González et al.
 (2020), $\phi$ was used as an observable proxy for the degree of symmetry in each system to determine whether or not the correction factor was needed.
-Larger-Einstein-radius systems with a predominantly low $\phi$ tend to deviate from a symmetric dark-matter halo and thus require the empirical correction.
+Larger-Einstein-radius systems with a predominantly low $\phi$ tend to deviate from a symmetric dark matter halo and thus require the empirical correction.
 For spherically symmetric systems with a large $\phi$, the measured $M_{\Sigma}(<\theta_{E})$ is taken to be fairly unbiased, and thus the correction factor is not needed.
-We obtained $\phi$ by determining the fraction of the perimeter of the Einstein aperture that was subtended by the the regions masking the lensed source arcs described in Section 2.4 in each system.
+We obtained $\phi$ by determining the fraction of the perimeter of the Einstein aperture that was subtended by the regions masking the lensed source arcs described in Section 2.4 in each system.
 Adopting the convention used by Remolina González et al.
@@
 If $\phi\geq 0.5$, it was not applied.
-Accounting for the empirical correction, the mean and standard deviation of the 1000 measurements of $M_{\Sigma}(<\theta_{E})$ were adopted as the enclosed total mass and its error respectively for each system.
+Accounting for the empirical correction, the mean and standard deviation of the 1000 measurements of $M_{\Sigma}(<\theta_{E})$ were adopted as the enclosed total mass and its error, respectively, for each system.
 3.2 Stellar Luminosity and Mass
@@
 A summary of all eight parameters used in fitting can be found in Table 1.
+While the number of free parameters used in fitting is greater than the number of photometric band constraints for each system, parameters such as age, metallicity, and dust are correlated with one another (e.g., Worthey 1994; Thomas et al.
+2017), and thus, SED fitting as done here remains sensible.
 Utilizing emcee (Foreman-Mackey et al., 2013) as implemented in Prospector with 84 walkers for a total of 6720 iterations, only the last 840 iterations for each of the 84 walkers (70,560 total parameter vectors) were taken as representing the posterior distribution for each of the five free parameters in each system in order to eliminate the significant burn-in sequence of the fitting process.
@@
 Observed-frame SEDs for the systems shown in Figure 2 after SED fitting are shown in Figure 3.
+Figure 4: A corner plot of the BCG redshift, enclosed total mass, and enclosed luminosity for the systems in this work plotted as black circles.
+Errors for each variable in each subplot are represented as red, blue, and green lines, respectively.
+A plane-of-best-fit is also shown in the top-right plot as well.
+This Figure is available as an animation.
+The animation depicts a rotating view of the plane and its associated constraints shown in the top right of the static Figure to emphasize its features in 3D.
+(An animation of this figure is available in the online article.)
 To derive luminosity, we generated new SEDs with a random sample of 1024 free-parameter vectors from the posterior distributions of the best-fit SED in each system.
-We integrated over the rest-frame wavelength interval of 3000Å to 7000Å, since this corresponds to the longest wavelength interval sampled by part of at least two bands of photometry across all BCG redshifts in this work (the lower bound of 3000Å gets redshifted out of the $\mathit{g}$ and into the $\mathit{r}$ band at high redshifts).
-This range also allows us to sample the rest-frame 4000Å break while recovering flux in longer wavelengths which are significantly more luminous in cluster galaxies.
-After integrating the SEDs, we converted flux to luminosity in solar luminosities using the appropriate luminosity distance and adopted the mean and standard deviation from the 1024 random parameter vectors as the luminosity and its error respectively for each system.
+We integrated over the rest-frame wavelength interval of 3000–7000Å, since this corresponds to the longest wavelength interval sampled by part of at least two bands of photometry across all BCG redshifts in this work (the lower bound of 3000Å gets redshifted out of the $\mathit{g}$ and into the $\mathit{r}$ band at high redshifts).
+This range also allows us to sample the rest-frame 4000Å break while recovering flux in longer wavelengths, which are significantly more luminous in cluster galaxies.
+After integrating the SEDs, we converted flux to luminosity in solar luminosities using the appropriate luminosity distance and adopted the mean and standard deviation from the 1024 random-parameter vectors as the luminosity and its error, respectively, for each system.
 4 Results
 As Figure 4 shows, the BCG redshift, enclosed total mass, and enclosed luminosity are correlated.
-Indeed, we note that the most massive clusters in this work also possess a large stellar mass and thus have a large luminosity.
-Using emcee, we constrained a plane-of-best-fit using a likelihood maximization estimation taking into account the error in data on all three axes to describe this correlation—where X = $z_{{}_{BCG}}$, Y = $\mathrm{log}_{10}[M_{\Sigma}(<\theta_{E})/M_{\odot}]$, and Z = $\mathrm{log}_{10}[L(<\theta_{E})/L_{\odot}]$—yielding:
-$Z=0.137_{-0.051}^{+0.051}X+0.522_{-0.023}^{+0.022}Y+4.207_{-0.290}^{+0.300}$ (4)
-This simple and analyic relationship derived from a sample of known strong lenses, considering only quantities derived from observations, may have utility in differentiating bona fide strong lenses from non-lenses in large samples of candidate lenses where the evidence of lensing is not nearly as robust as the sample of clusters analyzed here.
+By performing a linear regression scatter using emcee that accounts for errors in data on both axes (Hogg et al., 2010), we find the slope of the logarithmic correlation between the enclosed total mass and enclosed luminosity to be $0.528^{+0.022}_{-0.023}$ with an intrinsic scatter in luminosity of $0.115^{+0.008}_{-0.007}$ dex from the posterior distribution of the Markov Chain Monte Carlo (MCMC) fit.
+Indeed, we note that the most-massive clusters in this work also possess a large stellar mass and thus have a large luminosity.
+Also using emcee, we constrained a plane-of-best-fit taking into account the error in data on all three axes to describe this correlation—where X = $z_{{}_{BCG}}$, Y = $\mathrm{log}_{10}[M_{\Sigma}(<\theta_{E})/M_{\odot}]$, and Z = $\mathrm{log}_{10}[L(<\theta_{E})/L_{\odot}]$ with an intrinsic scatter on Z of $0.110^{+0.009}_{-0.008}$ dex—yielding:
+$Z=0.148^{+0.051}_{-0.052}X+0.529^{+0.023}_{-0.022}Y+4.121^{+0.272}_{-0.298}$ (4)
+The strong lensing candidates used to derive this relationship are, in our opinion, robust and unlikely to be non-lensing objects misclassified as bona fide lenses.
+Within the COOL-LAMPS candidate sample from which this subset derives, and indeed any other sample of strong lens candidates, there are a much larger number of less robust candidate systems.
+The identification of these systems as strong lenses may be more uncertain due to less obvious lensing geometry, and also—as is true in all flux-limited surveys—many of the notional lensed images will be only ambiguously detected.
+However, the simple and analytic relationship above, derived from a subsample of robust candidate strong lenses considering only quantities derived from the discovery survey observations, may have utility in differentiating bona fide strong lenses from non-lenses in the correspondingly larger samples of candidates where the evidence of lensing is not nearly as robust as the sample of candidates analyzed here.
 Moreover, this correlation offers a future opportunity to test simulations against observations—particularly those that aim to generate realistic galaxy populations in dense cluster environments.
-Figure 3: A 3D visualization of the BCG redshift, enclosed total mass, and enclosed luminosity for the systems in this work plotted as black circles.
-A plane-of-best-fit is also shown in black as well.
-Errors in data on each axis are represented as red, blue, and green lines respectively.
 A full accounting of all measured quantities in this work can be found in Table 2.
+Figure 5: The total mass enclosed within the Einstein aperture for each system on the $x$-axis plotted against both the enclosed stellar mass and the enclosed stellar-to-total mass fraction on the $y$-axes.
+Points in Panels I and III are color coded according to the luminosity of the system in units of solar luminosities from stellar light enclosed within the Einstein aperture as measured within the rest-frame wavelength interval of 3000–7000Å.
+Points in Panels II and IV only represent systems in which multiple cluster members fall within the Einstein aperture.
+Linear regression lines in all four Panels are drawn from the posterior distributions of their respective MCMC.
 5 Discussion
 5.1 Scaling Relations
-Besides the enclosed total mass and luminosity, we can also infer the observed enclosed stellar mass at the redshift of observation for each system using the stellar-mass posterior distribution from each SED constrained with Prospector.
-We obtained this by correcting the derived enclosed stellar mass formed in each system with the surviving mass fraction returned by Prospector.
-This is necessary because the amount of stellar mass that we see at the redshift of observation is less than the total stellar mass formed over the lifetime of the galaxy (e.g., Li et al.
-2017)—the latter being what Prospector parametrically fits.
+In addition to the enclosed total mass and luminosity, we can also infer the observed enclosed stellar mass at the redshift of observation for each system.
+The stellar-mass posterior distribution from each SED constrained with Prospector describes the total amount of stellar mass formed over the lifetime of each system.
+However, the amount of stellar mass that we see at the redshift of observation is less than the total stellar mass formed over the lifetime of the system (e.g., such as supernovae or outflows; Li et al.
+2017).
+To correct for this, Prospector also returns a surviving mass fraction at the conclusion of each fit that describes the percentage of the total stellar mass that remains in the system at the redshift of observation.
+By multiplying this percentage by the best-fit stellar mass formed, we obtain the stellar mass at the redshift of observation for each system.
 While we note that enclosed stellar mass is more complex to infer and subject to more significant systematics when compared to the more direct measures of enclosed total mass from strong-lensing geometry and luminosity, it nevertheless offers another point of comparison with those measurables.
 5.1.1 Total Mass—Stellar Mass
-Figure 4: The total mass enclosed within the Einstein aperture for each system on the x-axis plotted against both the enclosed stellar mass and the enclosed stellar-to-total mass fraction on the y-axes.
-Points in Panels I and III are color coded according to the luminosity of the system in units of solar luminosities from stellar light enclosed within the Einstein aperture as measured within the rest-frame wavelength interval of 3000Å to 7000Å.
-Points in Panels II and IV only represent systems in which multiple cluster members fall within the Einstein aperture.
-Linear regression lines in all four Panels are drawn from the posterior distributions of their respective MCMC.
 In Panel I of Figure 5, we see a clearly positive correlation between the enclosed total mass and the enclosed stellar mass for all systems.
-By performing a linear regression with emcee using a likelihood maximization estimation that accounts for errors in data on both axes, we find that the slope of this correlation is $0.443\pm 0.035$ from the posterior distribution of the MCMC fit.
+By performing a linear regression with emcee that accounts for errors in data on both axes, we find that the slope of this correlation is $0.500^{+0.029}_{-0.031}$ with an intrinsic scatter of $0.104^{+0.005}_{-0.003}$ dex from the posterior distribution of the MCMC fit.
 We colored the points according to their luminosity to highlight how the intrinsic differences between individual systems contribute to the overall intrinsic scatter.
 In Table 2 of Kravtsov et al.
-(2018), they found $M_{500}$ (White, 2001) and the stellar mass of solely the BCG to be correlated with a logarithmic slope of $0.33\pm 0.11$.
-In addition, Kravtsov et al.
-(2018) found $M_{500}$ and the stellar mass of the entire cluster to be correlated with a logarithmic slope of $0.59\pm 0.08$.
-The fact that the slope for this work falls between the two slopes quoted by Kravtsov et al.
-(2018) is consistent with how our methodology samples a physical regime between exclusively the BCG and that of the entire cluster.
-Panel II of Figure 4 further explores this correlation, with only those clusters where multiple cluster members were contained within the Einstein aperture are plotted.
-A linear regression for the multiple-galaxy systems gives a slope of $0.453\pm 0.049$, which is consistent with the slope using all systems.
-5.1.2 Total Mass — Stellar-to-Total Mass Fraction
-In Panel III of Figure 4, we plot the enclosed stellar-to-total mass fraction against the enclosed total mass and see a clearly negative correlation (again with the same colormapping as Panel I to emphasize intrinsic scatter).
-Also using emcee, we find that the slope of this correlation is $-0.563\pm 0.035$.
-Categorizing in Panel IV identically to Panel II, we find that the multiple-galaxy systems are correlated with a slope of $-0.561\pm 0.050$.
-We again find that the slope for the multiple-galaxy systems are statistically indistinguishable from the slope for the entire sample shown in Panel III.
-Andreon (2010) measured 52 clusters over a slightly larger mass range and found that the stellar mass fraction depended on halo mass with a slope of $-0.55\pm 0.08$ when measuring out to a physical extent of $R_{200}$.
-Both of the slopes in Panels III and IV are statistically similar to Andreon (2010), and while the linear fit derived by Andreon (2010) has a slightly larger normalization, this is to be expected because the cluster masses are constrained out to $R_{200}$.
+(2018), they found $M_{500}$ (White, 2001) and the total stellar mass of the cluster within $R_{500}$ to be correlated with a logarithmic slope of $0.59\pm 0.08$—in agreement with the slope we derive.
+Figure 6 shows a distribution of the physical scales within which mass and light was measured for this work and that of Kravtsov et al.
+(2018) and Gonzalez et al.
+(2013).
+It is striking that, despite differentiating between the small-scale apertures considered in this work as well as large-scale apertures with a physical distance scale up to 30 times larger, we measure an identical relationship across clusters in that the more-massive clusters are less efficient in converting baryons into stars as indicated by a sub-unity stellar-to-total mass slope.
+We caution against overinterpretation of this result given that, in both cases, the total and stellar masses are cylindrical masses projected through a 3D halo.
+Therefore, even a very small aperture samples some mass and galaxy light at large radii in projection.
+Nevertheless, this result suggests that cluster galaxies at a range of cluster-centric radii, at least those probed by the combination of our results and Kravtsov et al.
+(2018) and Gonzalez et al.
+(2013), follow similar efficiency dependencies as a function of total halo mass.
+In addition, since cluster galaxies at larger radii are located within larger apertures, yet retain the same slope between stellar and total mass, cluster galaxies in the core and the outskirts of the cluster (i.e.
+at a range of cluster-centric radii) must be similarly affected by the mass of the cluster at large in terms of the total baryon conversion efficiency.
+Panel II of Figure 5 further explores this regime, in which only those clusters where multiple cluster galaxies are contained within the Einstein aperture are plotted.
+We note that 103 of the clusters considered in this work met this criterion, and a linear regression for these systems gives a slope of $0.498^{+0.040}_{-0.039}$—identical to that of our entire sample.
+Figure 6: A log-space histogram of the fitted Einstein radii for this work scaled to physical units (kpc) shown in blue.
+Also shown is a log-space histogram of $r_{500}$ for 21 galaxy clusters analyzed in Kravtsov et al.
+(2018) and Gonzalez et al.
+(2013).
+Seeing as how the measured slopes agree across approximately two decades of cluster-centric radii suggests that whatever mechanism is responsible for suppressing baryon conversion efficiency is cluster wide and operating similarly regardless of the subpopulation of cluster member galaxies considered.
+What is unclear is whether this is predominantly a case of nature or nurture.
+Do the smaller halos that merge to form a massive cluster at later times have intrinsically lower baryon conversion efficiency, and that characteristic is inherited by the merged object (nature)?
+Or, is this suppression occurring at later times as a consequence of the halo growth through accretion and mergers (nurture)?
+That the stellar populations in cluster galaxies are much older than the assembled cluster-scale halo (e.g., Khullar et al.
+2022) suggests the former.
+5.1.2 Total Mass—Stellar-to-total-mass Fraction
+Andreon (2010) measured 52 clusters and found that the stellar mass fraction depended on $M_{200}$ with a slope of $-0.55\pm 0.08$.
+To compare directly with Andreon (2010), we plot the enclosed stellar-to-total mass fraction against the enclosed total mass in Panel III of Figure 5 and see a clearly negative correlation (again with the same color-mapping as Panel I to emphasize intrinsic scatter).
+Also using emcee, we find that this correlation has a slope of $-0.495^{+0.032}_{-0.033}$ and an intrinsic scatter of $0.103^{+0.005}_{-0.002}$ dex.
+Categorizing in Panel IV identically to Panel II, we find that the multiple-galaxy systems are correlated with a slope of $-0.497^{+0.043}_{-0.047}$ and intrinsic scatter of $0.105^{+0.007}_{-0.004}$ dex.
+We again find that the slope for the multiple-galaxy systems is statistically identical to the slope for the entire sample shown in Panel III.
+The slopes we derive in Panels III and IV are statistically indistinguishable from those of Andreon (2010), who measured out to the physical distance scale of $R_{200}$—systematically larger than $R_{500}$ used in Kravtsov et al.
+(2018) and Gonzalez et al.
+(2013) and much larger than this work.
+While a later paper by Andreon (2012) found a shallower slope of $-0.38\pm 0.07$ with a much smaller sample of only 12 galaxy clusters, the slope we derive with 177 clusters and that of Andreon (2010) supports the prescription of a slope of $\sim-0.49$ from Bode et al.
+(2009) to reproduce observed X-ray scaling relations without any supplemental feedback from nonstellar sources.
+Kravtsov et al.
+(2018) also noted that the steepness of the slope they derived, in agreement with our own, indicates weaker feedback than cosmological simulations (e.g., Kravtsov et al.
+2005; Martizzi et al.
+2014) suggest.
+While this conclusion was not explicitly probed in this work, the large difference between the aperture sizes we use and those of Andreon (2010) and Kravtsov et al.
+(2018) and the similarity of the derived slopes is suggestive of scale invariance with respect to feedback strength in addition to baryon conversion efficiency across a wide range of cluster-centric radii.
 5.2 Photometric Bias and Error
@@
 In addition, bluer galaxies at redshift $z>z_{{}_{BCG}}$ appear relatively fainter, which are also unlikely to bias our methodology of measuring cluster light.
-Interlopers such as foreground stars that appear in front of cluster galaxies are forced to be masked, but any underlying lens-galaxy flux is typically minimal due to the limited angular size of such stars.
+Interlopers such as foreground stars that appear in front of cluster galaxies are forced to be masked, but any underlying cluster galaxy flux is typically minimal due to the limited angular size of such stars.
 In general, we expect that all of the unmasked flux from cluster cores at the angular scales considered in this work originates from genuine cluster members.
@@
 While we must also consider associated galaxy populations in the immediate large-scale structure when considering galaxy clusters projected onto the sky—as such structures may also be projected onto the cluster core—those galaxies will also contribute to the lensing effect and will have their mass and light captured by the methodology presented here.
-We also note that our sample contains extremely few examples of systems which have a measured Einstein radius of $\theta_{E}\lessapprox 2^{\prime\prime}$.
+As shown in Figure 1 and quantified in Section 2.1, our sample becomes incomplete at a measured Einstein radius of $3.09^{\prime\prime}$—falling to 50% incomplete by $2.11^{\prime\prime}$.
 To some extent, this is a consequence of our chosen sample of cluster-scale lenses, which we would generally expect to have larger Einstein radii.
 However, we caution that there exists a large scatter between the total halo mass and Einstein radius (Fox et al., 2022).
-Systems whose arcs are separated from the central galaxy at small radii may also be located directly within the discernable BCG light.
+Systems whose arcs are separated from the central galaxy at small radii may also be located directly within the discernible BCG light.
 This would not only make them hard to visually notice, but it would also make them unsuitable for this analysis because the source light would be significantly blended together with the lensing cluster light.
-While we thus expect this work to undersample the small-Einstein-radius regime, it is not apparent that this significantly alters our findings.
-In addition, our use of relatively shallow ground-based imaging limits the $5\sigma$ point-source magnitude depth of objects in the $g$, $r$, and $z$-band filters to $\approx$ 24.7, 23.9, and 23.0 respectively in the average case of two exposures per filter (Dey et al., 2019).
-Systems with a lens redshift $z\gtrapprox 1$ are barely visible even in $\mathit{z}$-band LS DR9 imaging data, and they are also simply a rarer type of lensing system in the universe (Li et al., 2019).
-Both the data depth and intrinsic redshift distribution of lenses shape the redshift distribution in this work.
+While we thus expect and demonstrate that this work undersamples the small-Einstein-radius regime, it is not apparent that this significantly alters our findings.
+In addition, our use of relatively shallow ground-based imaging limits the $5\sigma$ point-source magnitude depth of objects in the $g$-, $r$-, and $z$-band filters to $\approx$ 24.7, 23.9, and 23.0, respectively in the average case of two exposures per filter (Dey et al., 2019).
+Systems with a lens redshift $z\gtrapprox 1$ are barely visible even in $\mathit{z}$-band LS DR9 imaging data, and they are also simply a rarer type of strong lens in the Universe (Li et al., 2019).
+Both the data depth and intrinsic redshift distribution of lenses shape the redshift distribution of the clusters in this work.
 5.3 Alternative Mass Measurement
 In constraining mass, we also considered a second method for measuring cluster-centric mass elucidated in Remolina González et al.
-(2021a) using the parametric lens-modelling software LENSTOOL (Jullo et al., 2007) to measure the enclosed total mass—as opposed to the simple evaluation of Equations (1), (2), and (3).
-For the first 35 clusters that were analyzed in this work, we generated single-halo lens models by taking the three coordinates in each lensed arc used to derive the Einstein radius in Section 2.1 as a multiple-image family with a single pseudo-isothermal elliptical mass distribution (PIEMD, Kassiola & Kovner 1993) locked to the center of the BCG.
-Mass estimates were then obtained from the resulting best-fit lens models according to the methodology described in Remolina González et al.
+(2021a) using the parametric lens-modeling software LENSTOOL (Jullo et al., 2007) to measure the enclosed total mass—as opposed to the simple evaluation of Equations (1), (2), and (3).
+For the first 35 clusters that were analyzed in this work, we generated single-halo lens models by taking the three coordinates in each lensed arc used to derive the Einstein radius in Section 2.1 as a multiple-image family with a single pseudo-isothermal elliptical mass distribution (Kassiola & Kovner, 1993) with a free ellipticity locked to the center of the BCG.
+Mass estimates were then obtained from the resulting best-fit lens models by integrating over the surface-mass-density image, a 2D distribution in the lens plane of the mass fit with LENSTOOL.
+In this image, each pixel value represents the number of solar masses contained within that pixel.
+A circular aperture centered on the BCG with area equal to the area enclosed by the best-fit tangential critical curves was used to solve for the LENSTOOL enclosed mass in keeping with Remolina González et al.
 (2021a).
-Figure 5: A comparison of the single-halo LENSTOOL-generated mass measurements ($M_{SHM}$) and the parametrically generated mass measurements ($M_{\Sigma}$) for the first 35 clusters analyzed in this work.
-The black line represents an ideal 1-1 correspondence, and the blue lines represent the uncertainty on the measured correspondence.
+Figure 7: A comparison of the single-halo LENSTOOL-generated mass measurements ($M_{SHM}$) and the functionally generated mass measurements ($M_{\Sigma}$) for the first 35 clusters analyzed in this work.
+The black line represents an ideal one-to-one correspondence, and the blue lines represent the uncertainty on the measured correspondence.
 As expected from Remolina González et al.
@@
 The two mass measurements for the 35 clusters studied here are correlated with a slope of $0.960\pm 0.036$ and intercept of $0.492\pm 0.476$ dex as shown in Figure 7.
-In other words, we find no reason to suggest a preference between one method over the other, and we abandoned the LENSTOOL approach in favor of the much faster parametric approach throughout the rest of the work.
-In theory, lens modelling has the potential to be the most accurate mass-constraining methodology.
+In other words, we find no reason to suggest a preference between one method over the other, and we abandoned the LENSTOOL approach in favor of the much faster functional approach throughout the rest of the work.
+The agreement between the free-ellipticity-model masses and the circularly symmetric functional masses also demonstrates the feasibility of using a functional mass model for the geometrically simple strong lenses considered in this work.
+Lens modeling has the potential to be the most accurate mass-constraining methodology in galaxy clusters.
 However, this requires high-spatial-resolution imaging data and ample time.
-The process of iteratively refining just one lens model may take upwards of an hour for even a single relatively simple system—which is untenable when the number of candidate systems reaches into the thousands.
-Moreover, the accuracy of measuring mass with lens modelling is sensitive to the choice of defining multiple-image families along the lensed source arcs in a way that simple circle fitting is not.
-These essential positional constraints are often unresolved even with the highest-quality ground-based imaging data.
+The process of iteratively refining just one lens model may take upward of an hour for even a single relatively simple system—which is untenable when the number of candidate systems reaches into the thousands.
+Moreover, the accuracy of measuring mass with lens modeling is sensitive to the choice of defining multiple-image families along the lensed source arcs in a way that simple circle fitting is not.
+The essential positional constraints that define these image families are often unresolved—even with the highest-quality ground-based imaging data.
 Conversely, the positional constraints that inform the Einstein radius ultimately used by Equations (1), (2), and (3) need not be extremely accurate with respect to exactly constraining image family positions.
-As long as the positional constraints are reasonably placed within the arcs, something which is resolved by ground-based imaging data, it is accurate to within the systematic uncertainty.
-For more information on lens-modelling software, related techniques, and applications of strong lensing, we refer the reader to the following non-exhaustive list of papers for further review: Elíasdóttir et al.
+As long as the positional constraints are reasonably placed within the arcs, something that is resolved by ground-based imaging data, it is accurate to within the systematic uncertainty.
+For more information on lens-modeling software, related techniques, and applications strong lensing, we refer the reader to the following non-exhaustive list of papers for further review: Elíasdóttir et al.
 (2007), Oguri (2010), Lefor et al.
@@
 6 Summary and Conclusions
-This paper presents measurements of mass and light in 177 strong-lensing systems at cluster and group mass scales.
-By directly obtaining the Einstein radius and BCG redshift for each system, we compute the Einstein-radius-enclosed total mass parametrically.
-We use the enclosed $g$, $r$, and $z$-band photometry with Prospector to parameterize the enclosed luminosity and stellar mass, and we obtain a planar relationship between the BCG redshift, enclosed total mass, and enclosed luminosity of the form $Z=0.137_{-0.051}^{+0.051}X+0.522_{-0.023}^{+0.022}Y+4.207_{-0.290}^{+0.300}$ where X = $z_{{}_{BCG}}$, Y = $\mathrm{log}_{10}[M_{\Sigma}(<\theta_{E})/M_{\odot}]$, and Z = $\mathrm{log}_{10}[L(<\theta_{E})/L_{\odot}]$.
-We find that the enclosed total mass and stellar mass for the entire sample of strong-lensing systems considered in this work are correlated with a logarithmic slope, normalization, and intrinsic stellar-mass scatter of $0.443\pm 0.035$, $5.958\pm 0.454$ dex, and $0.173\pm 0.013$ dex respectively.
-We also find that the enclosed total mass and stellar-to-total mass fraction are correlated with a logarithmic slope, normalization, and intrinsic stellar-to-total-mass-fraction scatter of $-0.563\pm 0.035$, $6.034\pm 0.457$ dex, and $0.161\pm 0.013$ dex respectively.
+This paper presents measurements of mass and light in 177 high-confidence cluster-scale strong lenses.
+By directly obtaining the BCG redshift and an estimate for the Einstein radius for each system, we compute the Einstein-radius-enclosed total mass functionally and independent of any light-traces-mass model.
+We use the enclosed $g$-, $r$-, and $z$-band LS DR9 photometry with Prospector to estimate the enclosed luminosity and stellar mass, and we obtain a planar relationship between the BCG redshift, enclosed total mass, and enclosed luminosity of the form $Z=0.148^{+0.051}_{-0.052}X+0.529^{+0.023}_{-0.022}Y+4.121^{+0.272}_{-0.298}$ with an intrinsic scatter on $Z$ of $0.110^{+0.009}_{-0.008}$ dex where $X=z_{{}_{BCG}}$, $Y=\mathrm{log}_{10}[M_{\Sigma}(<\theta_{E})/M_{\odot}]$, and $Z=\mathrm{log}_{10}[L(<\theta_{E})/L_{\odot}]$.
+We find that the enclosed total mass and stellar mass for the entire sample of strong lenses considered in this work are correlated with a logarithmic slope, normalization, and intrinsic stellar-mass scatter of $0.500^{+0.029}_{-0.031}$, $5.254^{+0.395}_{-0.391}$, $0.104^{+0.005}_{-0.003}$ dex, respectively.
+We also find that the enclosed total mass and stellar-to-total mass fraction are correlated with a logarithmic slope, normalization, and intrinsic stellar-to-total-mass-fraction scatter of $-0.495^{+0.032}_{-0.033}$, $5.24^{+0.409}_{-0.424}$, $0.103^{+0.005}_{-0.002}$ dex, respectively.
+We recover scaling relations identical to that of works that measure within cluster-scale apertures—despite the fact that we measure inside apertures with radii some 30–100 times smaller.
+This is suggestive of invariance in baryon conversion efficiency and feedback strength as a function of cluster-centric radii in galaxy clusters.
 Enormous volumes of imaging data from new space-based and ground-based surveys at deeper and sharper photometric limits are on the horizon.
-Probing these upcoming datasets with both visual and machine-learning driven searches for new strong lenses at a variety of scales is inevitable.
-The correlations shown in this work should have utility in filtering rankings of possible strong-lensing systems as well as offering another measure of cluster cores that can be compared to cosmological simulations.
+Probing these upcoming datasets with both visual and machine-learning-driven searches for new strong lenses at a variety of scales is inevitable.
+The correlations shown in this work should also have utility in ranking candidate strong lenses as well as offering another measure of cluster cores that can be compared to cosmological simulations.
 Acknowledgments
+The authors would like to thank the anonymous reviewer for providing detailed feedback, which significantly improved the manuscript.
+S.D.M.
+would also like to thank Allison Noble for insightful discussions surrounding the results and discussion Sections of this work.
 This work is supported by The College undergraduate program and the College Innovation Fund at the University of Chicago and the Department of Astronomy and Astrophysics at the University of Chicago.